### v1

# Accelerated Convergence of Frank–Wolfe Algorithms with Adaptive Bregman Step-Size Strategy

Abstract
We propose a Frank–Wolfe (FW) algorithm with an adaptive Bregman step-size strategy for smooth adaptable (also called: relatively smooth) (weakly-) convex functions. This means that the gradient of the objective function is not necessarily Lipschitz continuous, and we only require the smooth adaptable property. Compared to existing FW algorithms, our assumptions are less restrictive. We establish convergence guarantees in various settings, such as sublinear to linear convergence rates, depending on the assumptions for convex and nonconvex objective functions. Assuming that the objective function is weakly convex and satisfies the local quadratic growth condition, we provide both local sublinear and local linear convergence regarding the primal gap. We also propose a variant of the away-step FW algorithm using Bregman distances. We establish global accelerated (up to linear) convergence for convex optimization under the Hölder error bound condition and its local linear convergence for nonconvex optimization under the local quadratic growth condition over polytopes. Numerical experiments on nonnegative linear inverse problems, $\ell_{p}$ loss problems, phase retrieval, low-rank minimization, and nonnegative matrix factorization demonstrate that our proposed FW algorithms outperform existing methods.
1 Introduction
In this paper, we consider constrained optimization problems of the form
$\min_{x\in P}\quad f(x),$ (1.1)
where $f:\mathbb{R}^{n}\to(-\infty,+\infty]$ is a continuously differentiable function and $P\subset\mathbb{R}^{n}$ is a compact convex set. We are interested in both convex and nonconvex $f$ and assume that we have first-order oracle access to $f$, i.e., given $x\in\mathbb{R}^{n}$, we can compute $\nabla f(x)$. Usually, Problem (1.1) is solved by a variant of the projected gradient method (see, for example, [Nesterov2018-gj, Nocedal2006-sb]). However, even if $P$ is convex, it is not always possible to access the projection operator of $P$ or compute the projection efficiently. While interior point methods are a potential alternative approach to constrained optimization problems, their updates often require second-order information, which can be computationally expensive for large-scale problems. To alleviate the situation, Frank and Wolfe [Frank1956-cq] proposed the Frank–Wolfe (FW) algorithm (also known as the conditional gradient method), which is a projection-free first-order method. Instead of requiring access to a projection oracle, the FW algorithm requires only access to a so-called linear minimization oracle (LMO), which for a given linear function $a\in\mathbb{R}^{n}$ computes $y\in\operatorname*{argmin}_{v\in P}\langle a,v\rangle$. Linear minimization oracles are often much cheaper than projection oracles, as shown in [CP2021] (see also summary in [Braun2022-ew, Table 1.1]), so that, in practice, FW algorithms are often faster than projected gradient methods when the projection operation is not trivial. Additionally, FW algorithms tend to be numerically quite robust and stable due to the affine-invariance of the algorithm and can be also used, e.g., to provide theoretical guarantees for the approximate Carathéodory problem [cp2019approxCara].
The original FW algorithm is given in Algorithm 1.
Input: Initial point $x_{0}\in P$, objective function $f$, step-size strategy $\gamma_{t}\in[0,1]$
1
2for $t=0,\ldots$ do
3      $v_{t}\leftarrow\operatorname*{argmin}_{v\in P}\langle\nabla f(x_{t}),v\rangle$
4     $x_{t+1}\leftarrow(1-\gamma_{t})x_{t}+\gamma_{t}v_{t}$
Algorithm 1 Frank–Wolfe algorithm
Related Work
In this section, we briefly review the work most closely related to this paper. The FW algorithm was originally proposed by Frank and Wolfe [Frank1956-cq] and was independently rediscovered and extended by Levitin and Polyak [Levitin1966-sm] as the conditional gradient method; we will use these terms interchangeably. Canon and Cullum [Canon1968-sr] established an initial lower bound for the convergence rate of the FW algorithm, which was later improved by GuéLat and Marcotte  [GueLat1986-bg], which also provided the first analysis of the away-step FW algorithm by Wolfe [Wolfe1970-ax]. Jaggi [Jaggi2013-oz] provided a more detailed convergence analysis of the FW algorithm, establishing a new lower bound that demonstrated a trade-off between sparsity and error. Concurrently, Lan [lan2013complexity] examined the complexity of linear programming-based first-order methods, establishing a similar complexity lower bound.
The classical FW algorithm is known to zigzag when approaching the optimal face containing the optimal solution $x^{*}$. This behavior inspired Wolfe [Wolfe1970-ax] to propose the away-step FW algorithm in 1970 that shortcuts the zigzagging by removing atoms that slow down the iterate sequence. Lacoste-Julien and Jaggi [Lacoste-Julien2015-gb] showed the linear convergence of the away-step FW algorithm. Their analysis introduced a geometric constant, the pyramidal width, that measures the conditioning of the polytope $P$, representing the feasible region. Similar analyses based on this constant have been developed for other advanced variants, such as the pairwise FW algorithm [Lacoste-Julien2015-gb], Wolfe’s algorithm [Lacoste-Julien2015-gb], the blended conditional gradient method [BPTW2018], and blended pairwise conditional gradient methods and similar variants [CP2020boost, TTP2021]. Even Nesterov-style acceleration is possible under weak assumptions [CDP2019] building upon this analysis.
Pedregosa et al. [Pedregosa2020-ay] introduced a powerful adaptive step-size strategy that dynamically approximates the local Lipschitz continuity of the function $f$. They proved that this strategy is at least as effective as the traditional short-step strategy, and it was later enhanced by Pokutta [Pokutta2024-qu] to improve numerical stability. For nonconvex functions, Lacoste-Julien [Lacoste-Julien2016-mj] established sublinear convergence, and Maskan et al. [Maskan2025-oa] incorporated the difference of convex functions (DC) into FW algorithms. For a comprehensive review of FW algorithms, we refer the interested reader to the survey by Braun et al. [Braun2022-ew] and the brief introduction by Pokutta [Pokutta2024-qu].
Establishing the convergence rate of FW algorithms typically requires that $\nabla f$ is Lipschitz continuous, although exceptions exist, such as in the case of generalized self-concordant functions [CBP2021, SelfConcordantFW_2020]. Some first-order methods using the Bregman distance do not require the global Lipschitz continuity of $\nabla f$. Bolte et al. [Bolte2018-zt] proposed the Bregman proximal gradient algorithm and established its global convergence under the smooth adaptable property (see Definition 2.6). Hanzely et al. [Hanzely2021-sz] introduced accelerated Bregman proximal gradient algorithms. Takahashi et al. [Takahashi2022-ml] developed a Bregman proximal gradient algorithm for the DC optimization and its accelerated version. Subproblems of Bregman proximal gradient algorithms are often not solvable in closed form. Takahashi and Takeda [Takahashi2024-wz] used approximations to Bregman distances to solve subproblems in closed form. Rebegoldi et al. [Rebegoldi2018-pa] proposed an inexact version of the Bregman proximal gradient algorithm.
Contribution
Our contributions can be roughly summarized as follows:
Frank–Wolfe for Smooth Adaptable functions
We study Frank–Wolfe algorithms for smooth adaptable functions (also called: relatively smooth) functions. Some functions whose gradients are not Lipschitz continuous satisfy the smooth adaptable property with kernel generating distances (see Definition 2.1). The class of smooth adaptable functions appears in many applications, such as nonnegative linear inverse problems [Bauschke1997-vk, Takahashi2024-wz], $\ell_{p}$ loss problems [Kyng2015-uo, Maddison2021-nt] phase retrieval [Bolte2018-zt, Takahashi2022-ml], nonnegative matrix factorization (NMF) [Mukkamala2019-mk, Takahashi2024-cu], and blind deconvolution [Takahashi2023-uh]. While the gradient of any $\mathcal{C}^{2}$ function over compact sets is Lipschitz continuous, an $\ell_{p}$ loss function is not $\mathcal{C}^{2}$ when $21<p<$ 2, i.e., its gradient is not Lipschitz continuous over compact sets. Moreover, functions in these applications are weakly convex on compact sets. We establish sublinear to linear convergence rates for the respective variants, which are compatible and in line with existing results.
Adaptive Bregman Step-Size Strategy
We propose an adaptive Bregman step-size strategy and establish its convergence results for convex and nonconvex optimization. Table 1 summarizes our contributions for the various cases that we consider. While Vyguzov and Stonyakin [Vyguzov2024-tp] also proposed an adaptive FW algorithm with the Bregman distance similar to ours for convex optimization problems. They established a linear convergence rate assuming that the objective function is relatively strongly convex, the optimal solution belongs to $\operatorname{int}P$, and the angle condition holds; in contrast, we establish linear convergence under much weaker assumptions. Beyond that, we also establish sublinear convergence for convex optimization for the case of general relatively smooth functions. Assuming that $f$ is convex, satisfies the Hölder error bound condition, and the optimal solution $x^{*}\in\operatorname{int}P$, we also provide the accelerated (up to linear) convergence, however, in a different way than [Vyguzov2024-tp] with fewer assumptions. For nonconvex optimization, assuming that $f$ is weakly convex and satisfies the local quadratic growth condition, we establish local linear convergence for our step-size strategy. These results include a special case that $\nabla f$ is Lipschitz continuous, i.e., $L$-smoothness. Without weak convexity, we prove the global convergence rate of the FW algorithm to a stationary point of (1.1), similar to [Lacoste-Julien2016-mj].
Away-Step FW Algorithm
We propose a variant of the away-step FW algorithm with the Bregman distance and establish its linear convergence under the Hölder error bound condition for convex optimization and under the local quadratic growth condition for weakly convex optimization.
Numerical Experiments
In numerical experiments, we have applied the FW algorithm with the adaptive Bregman step-size strategy to nonnegative linear systems [Bauschke2017-hg], $\ell_{p}$ loss problems [Kyng2015-uo, Maddison2021-nt], phase retrieval [Patterson1934-sg, Patterson1944-nm], low-rank minimization [Dragomir2021-rv], and NMF [Gillis2020-dh]. Bauschke et al. [Bauschke2017-hg] showed applications to linear inverse problems to solve nonnegative linear systems. Maddison et al. [Maddison2021-nt] considered $\ell_{p}$ loss problems, and Kyng et al. [Kyng2015-uo] showed its application as Lipschitz learning on graphs. Bolte et al. [Bolte2018-zt] and Takahashi et al. [Takahashi2022-ml] applied Bregman proximal algorithms to phase retrieval, Dragomir et al. [Dragomir2021-rv] applied them to low-rank minimization, and Mukkamala and Ochs [Mukkamala2019-mk] and Takahashi et al. [Takahashi2024-cu] applied them to NMF. Throughout these applications, we compare our proposed algorithms with existing FW algorithms and the mirror descent algorithm [Nemirovskij1983-sm] as well as away-step FW algorithms (see e.g. [Braun2022-ew, GueLat1986-bg, Lacoste-Julien2015-gb, Wolfe1970-ax]).
We would also like to stress that while we formulate some of the results for the case that $x^{*}\in\operatorname{int}P$, the results can be extended to the case that $x^{*}\in\operatorname{ri}P$, i.e., the relative interior of $P$, which we did not do for the sake of clarity. Basically, the analysis is performed in the affine space spanned by the optimal face of $P$ in that case; the interested reader is referred to [Braun2022-ew] for details on how to extend the results.
Table 1: Convergence rates for the Frank–Wolfe algorithm. For nonconvex functions, convergence is measured using the Frank-Wolfe gap, $\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle\leq\epsilon$, instead of the primal gap, $f(x_{t})-f^{*}\leq\epsilon$. Weak convexity can be replaced by twice continuous differentiability (see Proposition 2.10). Convergence rates for the weakly convex optimization locally hold. AFW, HEB, and quad. denote the away-step Frank–Wolfe algorithm, the Hölder error bound condition with $q\geq 1$, and the quadratic growth condition, respectively.
FW Assumptions Convergence rate
$f$ convexity $f$ growth $x^{*}\in\operatorname{int}P$ $P$ polytope $L$-smooth $L$-smad ($\nu\in(0,1]$)
any convex $\mathcal{O}(\epsilon^{-1})$ [Jaggi2013-oz] $\mathcal{O}(\epsilon^{-1/\nu})$ ([Vyguzov2024-tp], Thm. 4.2)
[…]
any weak quad. [^3] We assume that $f$ is local quadratic growth. $\mathcal{O}(\epsilon^{-1})$ $\mathcal{O}(\epsilon^{-1/\nu})$ (Thm. 5.5)
[…]
any [^5] Its rate is established for the Frank–Wolfe gap. nonconvex $\mathcal{O}(\epsilon^{-2})$ [Lacoste-Julien2016-mj] $\mathcal{O}(\epsilon^{-1-1/\nu})$ (Thm. 5.1)
Outline
The structure of this paper is as follows. Section 2 introduces essential notations, including Bregman distances, the smooth adaptable property, weak convexity, growth conditions, and scaling inequalities, which are crucial for the convergence analyses that come later. In Section 3, we present the FW algorithm with an adaptive Bregman step-size strategy and the away-step FW algorithm utilizing the Bregman distance. Section 4 details the convergence results for convex optimization, demonstrating both sublinear and linear convergence rates. Section 5 addresses convergence results for the nonconvex case, showing that our algorithm achieves global convergence to a stationary point of (1.1). Under the assumption that $f$ is weakly convex and local quadratic growth, we establish local linear convergence to a minimizer of (1.1), indicating that the proposed algorithms converge to a minimizer when the initial point is sufficiently close. Section 6 presents applications to nonnegative linear inverse problems, phase retrieval, low-rank minimization, and nonnegative matrix factorization (NMF).
2 Preliminaries
[…]
Let $\operatorname{int}{C}$, $\operatorname{ri}{C}$, and $\operatorname{cl}{C}$ be the interior, the relative interior, and the closure of a set $C\subset\mathbb{R}^{n}$, respectively. The set $\operatorname{conv}C$ denotes the convex hull of $C$, and the set $\operatorname{Vert}C$ denotes the set of vertices of $C$. We also define the distance from a point $x\in\mathbb{R}^{n}$ to $C$ by $\operatorname{dist}(x,C):=\inf_{y\in C}\|x-y\|$.
For an extended-real-valued function $f:\mathbb{R}^{n}\to[-\infty,+\infty]$, the effective domain of $f$ is defined by $\operatorname{dom}f:=\{x\in\mathbb{R}^{n}\,|\,f(x)<+\infty\}$. The function $f$ is said to be proper if $f(x)>-\infty$ for all $x\in\mathbb{R}^{n}$ and $\operatorname{dom}f\neq\emptyset$. Let $[f\leq\zeta]:=\{x\in\mathbb{R}^{n}\,|\,f(x)\leq\zeta\}$ be a $\zeta$-sublevel set of $f$ for some $\zeta\in\mathbb{R}$. Let $\mathcal{C}^{k}$ be the class of $k$-times continuously differentiable functions for $k\geq 0$. The sign function $\operatorname{sgn}(x)$ is defined by $\operatorname{sgn}(x)=-1$ for $x<0$, $\operatorname{sgn}(x)=0$ for $x=0$, and $\operatorname{sgn}(x)=1$ for $x>0$.
2.1 Bregman Distances
Let $C$ be a nonempty open convex subset of $\mathbb{R}^{n}$.
Definition 2.1 (Kernel Generating Distance [Bolte2018-zt, Definition 2.1]).
A function $\phi:\mathbb{R}^{n}\to(-\infty,+\infty]$ is called a kernel generating distance associated with $C$ if it satisfies the following conditions:
  1. (i)
    $\phi$ is proper, lower semicontinuous, and convex, with $\operatorname{dom}\phi\subset\operatorname{cl}C$ and $\operatorname{dom}\partial\phi=C$.
  2. (ii)
    $\phi$ is $\mathcal{C}^{1}$ on $\operatorname{int}\operatorname{dom}\phi\equiv C$.
We denote the class of kernel generating distances associated with $C$ by $\mathcal{G}(C)$.
Definition 2.2 (Bregman Distance [Bregman1967-rf]).
Given a kernel generating distance $\phi\in\mathcal{G}(C)$, a Bregman distance $D_{\phi}:\operatorname{dom}\phi\times\operatorname{int}\operatorname{dom}\phi\to\mathbb{R}_{+}$ associated with $\phi$ is defined by
$D_{\phi}(x,y)=\phi(x)-\phi(y)-\langle\nabla\phi(y),x-y\rangle.$ (2.1)
The Bregman distance $D_{\phi}(x,y)$ measures the proximity between $x\in\operatorname{dom}\phi$ and $y\in\operatorname{int}\operatorname{dom}\phi$. Indeed, since $\phi$ is convex, it holds that $D_{\phi}(x,y)\geq 0$ for all $x\in\operatorname{dom}\phi$ and $y\in\operatorname{int}\operatorname{dom}\phi$. Moreover, when $\phi$ is strictly convex, $D_{\phi}(x,y)=0$ holds if and only if $x=y$. However, the Bregman distance is not always symmetric and does not have to satisfy the triangle inequality. The Bregman distance is also called the Bregman divergence.
Example 2.3.
Well-known choices for $\phi$ and $D_{\phi}$ are listed below; for more examples, see, e.g., [Bauschke2017-hg, Example 1], [Bauschke1997-vk, Section 6], and [Dhillon2008-zz, Table 2.1].
  1. (i)
    Let $\phi(x)=\frac{1}{2}\|x\|^{2}$ and $\operatorname{dom}\phi=\mathbb{R}^{n}$. Then, the Bregman distance corresponds to the squared Euclidean distance, i.e., $D_{\phi}(x,y)=\frac{1}{2}\|x-y\|^{2}$.
  2. (ii)
    The Boltzmann–Shannon entropy $\phi(x)=\sum_{i=1}^{n}x_{i}\log x_{i}$ with $00\log 0=$ 0 and $\operatorname{dom}\phi=\mathbb{R}_{+}^{n}$. Then, $D_{\phi}(x,y)=\sum_{i=1}^{n}(x_{i}\log\frac{x_{i}}{y_{i}}-x_{i}+y_{i})$ is called the Kullback–Leibler divergence [Kullback1951-yv].
  3. (iii)
    The Burg entropy $\phi(x)=-\sum_{i=1}^{n}\log x_{i}$ and $\operatorname{dom}\phi=\mathbb{R}_{++}^{n}$. Then, $D_{\phi}(x,y)=\sum_{i=1}^{n}(\frac{x_{i}}{y_{i}}-\log\frac{x_{i}}{y_{i}}-1)$ is called the Itakura–Saito divergence [Itakura1968-en].
  4. (iv)
    Let $\phi(x)=\frac{1}{4}\|x\|^{4}+\frac{1}{2}\|x\|^{2}$ and $\operatorname{dom}\phi=\mathbb{R}^{n}$. The Bregman distance $D_{\phi}$ is used in phase retrieval [Bolte2018-zt], low-rank minimization [Dragomir2021-rv], NMF [Mukkamala2019-mk], and blind deconvolution [Takahashi2023-uh].
We recall the triangle scaling property for Bregman distances from Hanzely et al. [Hanzely2021-sz, Section 2], where several properties of the triangle scaling property are also shown.
Definition 2.4 (Triangle Scaling Property [Hanzely2021-sz, Definition 2]).
Given a kernel generating distance $\phi\in\mathcal{G}(C)$, the Bregman distance $D_{\phi}$ has the triangle scaling property if there exists a constant $\nu>0$ such that, for all $x,y,z\in\operatorname{int}\operatorname{dom}\phi$ and all $\gamma\in[0,1]$, it holds that
$D_{\phi}((1-\gamma)x+\gamma y,(1-\gamma)x+\gamma z)\leq\gamma^{\nu}D_{\phi}(y,z).$ (2.2)
Now, substituting $z\leftarrow x$ on the left-hand side of (2.2), we obtain
$D_{\phi}((1-\gamma)x+\gamma y,x)$ $=\phi((1-\gamma)x+\gamma y)-\phi(x)-\gamma\langle\nabla\phi(x),y-x\rangle$
$\leq(1-\gamma)\phi(x)+\gamma\phi(y)-\phi(x)-\gamma\langle\nabla\phi(x),y-x\rangle=\gamma D_{\phi}(y,x),$
where the inequality holds because of the convexity of $\phi$. Therefore, there exists $\nu>0$ such that $D_{\phi}((1-\gamma)x+\gamma y,x)\leq\gamma^{\nu}D_{\phi}(y,x)$ holds for all $x,y\in\operatorname{int}\operatorname{dom}\phi$ and all $\gamma\in[0,1]$.
We will now show that a stronger version can be obtained if $\phi$ is strictly convex, so that we can rephrase $\nu$ with $1+\nu$ where $\nu\in(0,1]$, i.e., the right-hand side is superlinear, which will be crucial for the convergence analysis later.
Lemma 2.5.
Given a kernel generating distance $\phi\in\mathcal{G}(C)$, if $\phi$ is strictly convex, then there exists $\nu\in(0,1]$ such that, for all $x,y\in\operatorname{int}\operatorname{dom}\phi$ and all $\gamma\in[0,1]$, it holds that
$D_{\phi}((1-\gamma)x+\gamma y,x)\leq\gamma^{1+\nu}D_{\phi}(y,x).$ (2.3)
Proof.
Obviously, (2.3) holds if $y=x$ or $\gamma\in\{0,1\}$. In what follows, we assume $y\neq x$ and $10<\gamma<$ 1. Let $g(\nu):=\gamma^{1+\nu}D_{\phi}(y,x)-D_{\phi}((1-\gamma)x+\gamma y,x)$ for all $x,y\in\operatorname{int}\operatorname{dom}\phi$, $y\neq x$ and all $\gamma\in(0,1)$. Using $D_{\phi}(y,x)>0$ due to the strict convexity of $\phi$, we have, for any $\nu\geq 0$,
$g^{\prime}(\nu)=\gamma^{1+\nu}D_{\phi}(y,x)\log\gamma<0,$
which implies $g$ monotonically decreases. In addition, it holds that
$g(0)$ $=\gamma D_{\phi}(y,x)-D_{\phi}((1-\gamma)x+\gamma y,x)$
$=\gamma(\phi(y)-\phi(x)-\langle\nabla\phi(x),y-x\rangle)-\phi((1-\gamma)x+\gamma y)+\phi(x)+\gamma\langle\nabla\phi(x),y-x\rangle$
$=(1-\gamma)\phi(x)+\gamma\phi(y)-\phi((1-\gamma)x+\gamma y)>0,$
where the last inequality holds because $\phi$ is strictly convex, i.e., $\phi((1-\gamma)x+\gamma y)<(1-\gamma)\phi(x)+\gamma\phi(y)$. Therefore, there exists $\nu>0$ such that $g(\nu)\geq 0$ by the intermediate value theorem. ∎
Because $\phi$ is quadratic when $D_{\phi}$ is symmetric [Bauschke2001-ry, Lemma 3.16], it always holds that $D_{\phi}((1-\gamma)x+\gamma y,x)=\gamma^{1+\nu}D_{\phi}(y,x)$ with $\nu=1$.
In the remainder of the paper, if not stated otherwise, we will assume that $\phi$ is strictly convex and as such for a given $\phi$ and $C$ there exists a $\nu\in(0,1]$ such that $D_{\phi}((1-\gamma)x+\gamma y,x)\leq\gamma^{1+\nu}D_{\phi}(y,x)$ holds for all $x,y\in\operatorname{int}\operatorname{dom}\phi$ and all $\gamma\in[0,1]$.
2.2 Smooth Adaptable Property
We will now recall the smooth adaptable property, which is a generalization of $L$-smoothness and was first introduced by [Bauschke2017-hg]. The smooth adaptable property is also called relative smoothness [Lu2018-ii].
Definition 2.6 ($L$-smooth Adaptable Property [Bolte2018-zt, Definition 2.2]).
Consider a pair of functions $(f,\phi)$ satisfying the following conditions:
  1. (i)
    $\phi\in\mathcal{G}(C)$,
  2. (ii)
    $f:\mathbb{R}^{n}\to(-\infty,+\infty]$ is proper and lower semicontinuous with $\operatorname{dom}\phi\subset\operatorname{dom}f$, which is $\mathcal{C}^{1}$ on $C\equiv\operatorname{int}\operatorname{dom}\phi$.
The pair $(f,\phi)$ is said to be $L$-smooth adaptable (for short: $L$-smad) on $C$ if there exists $L>0$ such that $L\phi-f$ and $L\phi+f$ are convex on $C$.
The convexity of $L\phi-f$ and $L\phi+f$ plays a central role in developing and analyzing algorithms, and the smooth adaptable property implies the extended descent lemma.
Lemma 2.7 (Extended Descent Lemma [Bolte2018-zt, Lemma 2.1]).
The pair of functions $(f,\phi)$ is $L$-smad on $C$ if and only if for all $x,y\in\operatorname{int}\operatorname{dom}\phi$,
$|f(x)-f(y)-\langle\nabla f(y),x-y\rangle|\leq LD_{\phi}(x,y).$
From this, it can be seen that the $L$-smad property for $(f,\phi)$ provides upper and lower approximations for $f$ majorized by $\phi$ with $L>0$. In addition, if $\phi(x)=\frac{1}{2}\|x\|^{2}$ on $\operatorname{int}\operatorname{dom}\phi=\mathbb{R}^{n}$, Lemma 2.7 corresponds to the classical descent lemma. While the $L$-smad property might seem unfamiliar at first, it is a natural generalization of $L$-smoothness, and examples of functions $f$ and $\phi$ satisfying the $L$-smad property are given, e.g., in [Bauschke2017-hg, Lemmas 7 and 8][Bolte2018-zt, Lemma 5.1], [Dragomir2021-rv, Propositions 2.1 and 2.3], [Mukkamala2019-mk, Proposition 2.1], [Takahashi2024-wz, Proposition 24], [Takahashi2023-uh, Theorem 1], and [Takahashi2024-cu, Theorem 15].
The extended descent lemma immediately implies primal progress of FW algorithms under the $L$-smad property as a straightforward generalization of the $L$-smooth case.
Lemma 2.8 (Primal progress from the smooth adaptable property).
Let the pair of functions $(f,\phi)$ be $L$-smad with a strictly convex function $\phi$ and let $x_{+}=(1-\gamma)x+\gamma v$ with $x,v\in\operatorname{int}\operatorname{dom}\phi$ and $\gamma\in[0,1]$. Then it holds:
$f(x)-f(x_{+})\geq\gamma\langle\nabla f(x),x-v\rangle-L\gamma^{1+\nu}D_{\phi}(v,x).$
Proof.
Using Lemma 2.7 and substituting $x_{+}$ for $x$ and $x$ for $y$, we have
$f(x)-f(x_{+})$ $\geq\gamma\langle\nabla f(x),x-v\rangle-LD_{\phi}(x_{+},x).$
It holds that $D_{\phi}(x_{+},x)\leq\gamma^{1+\nu}D_{\phi}(v,x)$ for some $\nu\in(0,1]$ by Lemma 2.5. Therefore, this provides the desired inequality. ∎
Often the point $v\in P\subset\operatorname{int}\operatorname{dom}\phi$ in Lemma 2.8 is chosen as a Frank–Wolfe vertex, i.e.,
$v\in\operatorname*{argmax}_{u\in P}\langle\nabla f(x),x-u\rangle,$
but other choices, e.g., those arising from away-directions from the away-step FW method, are also possible, as we will see in Section 3.2.
2.3 Weakly Convex Functions
A function $f$ is called weakly convex if it becomes convex upon adding a quadratic perturbation.
Definition 2.9 (Weakly Convex Function [Davis2018-bb, Nurminskii1975-vp]).
A function $f:\mathbb{R}^{n}\to(-\infty,+\infty]$ is said to be $\rho$-weakly convex for some $\rho>0$ if the function $f+\frac{\rho}{2}\|\cdot\|^{2}$ is convex.
Obviously, $f$ is $\rho$-weakly convex and differentiable if and only if, for all $x,y\in\operatorname{dom}f$,
$f(y)-f(x)\geq\langle\nabla f(x),y-x\rangle-\frac{\rho}{2}\|y-x\|^{2}.$
For examples of such $f$, see [Davis2018-bb, Examples 3.1 and 3.2]. Moreover, it is easy to see that any $\mathcal{C}^{2}$ function defined on a compact set is weakly convex [Vial1983-oo, Proposition 4.11] and [Wu2007-dt], the rationale being that we can simply shift the function by the quadratic belonging to its smallest eigenvalue. Since the proof of this fact is omitted therein, we provide the argument for the sake of completeness below.
Proposition 2.10.
Let $P\subset\mathbb{R}^{n}$ be a compact set and $f$ be a proper and $\mathcal{C}^{2}$ function on $P$. Then, $f$ is weakly convex on $P$.
Proof.
Let $g=\lambda_{\min}(\nabla^{2}f(\cdot))$. Obviously, $g$ is continuous because $\lambda_{\min}(\cdot)$ and $\nabla^{2}f(\cdot)$ are continuous. Therefore, there exists the minimum value of $g$ on $P$ due to the continuity of $g$ and the compactness of $P$. For $\rho=|\min_{x\in P}g(x)|$, we have $\nabla^{2}f(x)+\rho I\succeq O$ for any $x\in P$, which implies $f+\frac{\rho}{2}\|\cdot\|^{2}$ is convex on $P$, i.e., $f$ is $\rho$-weakly convex on $P$. ∎
Minimizing a nonconvex $\mathcal{C}^{2}$-function over a compact set $P$ is equivalent to minimizing a weakly convex function over $P$. However, it is important to note that when $\rho$ becomes large, the assumptions of algorithms may not be satisfied (see also Theorems 5.5 and 5.9).
2.4 Growth Conditions
In [garber2015faster] (see also [Garber2020-tr]) the linear convergence of the FW algorithm for convex optimization under the quadratic growth condition, which is a weaker assumption than assuming strong convexity was established over strongly convex sets and later generalized to uniformly convex sets in [KAP2020] as well as to conditions weaker than quadratic growth in [KDP2018, KDP2018jour]. Local variants of these notions, as necessary, e.g., for the nonconvex case, have been studied in [KAP2021] in the context of Frank–Wolfe methods.
Definition 2.11 (Hölder Error Bound [Braun2022-ew, Definition 3.25], [alex17sharp, Definition 1.2] and Quadratic Growth Conditions [garber2015faster, Garber2020-tr, Liao2024-vd]).
Let $f:\mathbb{R}^{n}\to(-\infty,+\infty]$ be a proper lower semicontinuous function and $P\subset\mathbb{R}^{n}$ be a compact convex set. Let $\mathcal{X}^{*}\neq\emptyset$ be the set of optimal solutions, i.e., $\mathcal{X}^{*}:=\operatorname*{argmin}_{x\in P}f(x)$, and let $f^{*}=\min_{x\in P}f(x)$ and $\zeta>0$. The function $f$ satisfies the Hölder error bound condition on $P$ if there exists constants $q\geq 1$ and $\mu>0$ such that, for any $x\in P$,
$\operatorname{dist}(x,\mathcal{X}^{*})^{q}\leq\frac{q}{\mu}(f(x)-f^{*}).$ (2.4)
Particularly, when $q=2$, it is called the $\mu$-quadratic growth condition. Moreover, the function $f$ is said to be local $\mu$-quadratic growth (with $\zeta$) if there exists a constant $\mu>0$ such that, for any $x\in[f\leq f^{*}+\zeta]\cap P$, (2.4) holds with $q=2$.
The Hölder error bound condition shows sharpness bounds on the primal gap (see, e.g., [bolte17, alex17sharp]), which has also been extensively analyzed for Frank–Wolfe methods in [KDP2018, KDP2018jour]. The convergence analysis for nonconvex optimization under the quadratic growth or Hölder error bound condition was established in [Davis2024-rm, Davis2018-bb, Davis2024-og]. We distinguish between the global condition and the local condition of Definition 2.11. We employ the Hölder error bound condition for convex optimization while we use the local quadratic growth for nonconvex optimization.
Next, we show in the following lemma that the Hölder error bound condition immediately provides a bound on the primal optimality gap, which will be useful for establishing convergence rates of the proposed algorithms as well as the Łojasiewicz inequality [Bolte2007-xr, Lojasiewicz1963-sm, Lojasiewicz1965-ks], provided $f$ is convex.
Lemma 2.12 (Primal gap bound from Hölder error bound).
Let $f$ be a convex function and satisfy the Hölder error bound condition with $q\geq 1$ and $\mu>0$. Let $x^{*}$ be the unique minimizer of $f$ over $P$. Then, the following argument holds in general, for all $x\in P$:
$f(x)-f^{*}\leq\left(\frac{q}{\mu}\right)^{\frac{1}{q-1}}\left(\frac{\langle\nabla f(x),x-x^{*}\rangle}{\|x-x^{*}\|}\right)^{\frac{q}{q-1}},$ (2.5)
or equivalently,
$\left(\frac{\mu}{q}\right)^{1/q}(f(x)-f^{*})^{1-1/q}\leq\frac{\langle\nabla f(x),x-x^{*}\rangle}{\|x-x^{*}\|}.$
Proof.
By first applying convexity and then the Hölder error bound condition for any $x^{*}\in\mathcal{X}^{*}$ with $f^{*}=f(x^{*})$ it holds:
$f(x)-f^{*}$ $\leq\langle\nabla f(x),x-x^{*}\rangle$
$=\frac{\langle\nabla f(x),x-x^{*}\rangle}{\|x-x^{*}\|}\|x-x^{*}\|$
$\leq\frac{\langle\nabla f(x),x-x^{*}\rangle}{\|x-x^{*}\|}\left(\frac{q}{\mu}(f(x)-f^{*})\right)^{1/q},$
which implies
$\left(\frac{\mu}{q}\right)^{1/q}(f(x)-f^{*})^{1-1/q}\leq\frac{\langle\nabla f(x),x-x^{*}\rangle}{\|x-x^{*}\|},$
or equivalently
$f(x)-f^{*}\leq\left(\frac{q}{\mu}\right)^{\frac{1}{q-1}}\left(\frac{\langle\nabla f(x),x-x^{*}\rangle}{\|x-x^{*}\|}\right)^{\frac{q}{q-1}}.$
Remark 2.13.
With the remark from above, we can immediately relate the Hölder error bound condition to the Łojasiewicz inequality [Bolte2007-xr, Lojasiewicz1963-sm, Lojasiewicz1965-ks]; the Łojasiewicz inequality is used to establish convergence analysis [Bolte2007-xr]. To this end, using (2.5), we estimate $\frac{\langle\nabla f(x),x-x^{*}\rangle}{\|x-x^{*}\|}\leq\frac{\|\nabla f(x)\|\|x-x^{*}\|}{\|x-x^{*}\|}=\|\nabla f(x)\|$ to then obtain the weaker condition:
$f(x)-f^{*}\leq\left(\frac{q}{\mu}\right)^{\frac{1}{q-1}}\|\nabla f(x)\|^{\frac{q}{q-1}}=c^{\frac{1}{\theta}}\|\nabla f(x)\|^{\frac{1}{\theta}},$ (2.6)
where $c=\left(\frac{q}{\mu}\right)^{\frac{1}{q}}$ and $\theta=\frac{q-1}{q}$. Inequality (2.6) is called the $c$-Łojasiewicz inequality with $\theta\in[0,1)$. If $\mathcal{X}^{*}\subseteq\operatorname{ri}P$, then the two conditions are equivalent. However, if the optimal solution(s) are on the boundary of $P$ as is not infrequently the case, then the two conditions are not equivalent as $\|\nabla f(x)\|$ might not vanish for $x\in\mathcal{X}^{*}$, whereas $\langle\nabla f(x),x-x^{*}\rangle$ does, i.e., the Hölder error bound condition is tighter than the one induced by the Łojasiewicz inequality.
The next lemma shows that we can also obtain primal gap bounds in the weakly convex case together with (local) quadratic growth. Its proof can be found in Appendix A.1.
Lemma 2.14 (Primal gap bound from the quadratic growth).
Let $f$ be a $\rho$-weakly convex function that satisfies the local $\mu$-quadratic growth condition such that $\rho\leq\mu$. Let $x^{*}$ be the unique minimizer of $f$ over $P$ and let $\zeta>0$. Then, the following holds: for all $x\in[f\leq f^{*}+\zeta]\cap P$,
$f(x)-f^{*}\leq\frac{2\mu}{(\mu-\rho)^{2}}\left(\frac{\langle\nabla f(x),x-x^{*}\rangle}{\|x-x^{*}\|}\right)^{2},$ (2.7)
or equivalently,
$\left(\frac{\mu}{2}\right)^{1/2}\left(1-\frac{\rho}{\mu}\right)(f(x)-f^{*})^{1/2}\leq\frac{\langle\nabla f(x),x-x^{*}\rangle}{\|x-x^{*}\|}.$
Remark 2.15.
In the same vein as the discussion in Remark 2.13, we can immediately relate the local $\mu$-quadratic growth condition with $\zeta>0$ to the local Polyak–Łojasiewicz (PL) inequality. Using (2.7), for all $x\in[f\leq f^{*}+\zeta]\cap P$, we have
$f(x)-f^{*}\leq\frac{2\mu}{(\mu-\rho)^{2}}\|\nabla f(x)\|^{2},$ (2.8)
which is equivalent to the PL inequality [Polyak1963-dr], also called the gradient dominance property [Combettes2020-jj], with $c=\frac{\sqrt{2\mu}}{\mu-\rho}$. The PL inequality is also equivalent to the Łojasiewicz inequality with $\theta=\frac{1}{2}$. Strongly convex functions satisfy the PL inequality [Braun2022-ew, Lemma 2.13]. Note that we will not have the primal gap bound under the Hölder error bound condition and weak convexity because an inequality like (A.1) does not hold from the Hölder error bound condition.
2.5 Scaling Inequalities, Geometric Hölder Error Bounds, and Contractions
In the following, we will now bring things together to derive tools that will be helpful in establishing convergence rates.
Scaling inequalities
Scaling inequalities are a key tool in establishing convergence rates of FW algorithms. We will introduce two such inequalities that we will use in the following. The first scaling inequality is useful for analyzing the case where the optimal solution lies in the relative interior of the feasible region. While its formulation in [Braun2022-ew, Proposition 2.16] required $L$-smoothness of $f$, it is actually not used in the proof, and the results hold more broadly. We restate it here for the sake of completeness.
Proposition 2.16 (Scaling inequality for inner optima from convexity [Braun2022-ew, Proposition 2.16]).
Let $P\subset\mathbb{R}^{n}$ be a nonempty compact convex set. Let $f:\mathbb{R}^{n}\to(-\infty,+\infty]$ be $\mathcal{C}^{1}$ and convex on $P$. If there exists $r>0$ so that $B(x^{*},r)\subset P$ for a minimizer $x^{*}$ of $f$, then for all $x\in P$, we have
$\langle\nabla f(x),x-v\rangle\geq r\|\nabla f(x)\|\geq\frac{r\langle\nabla f(x),x-x^{*}\rangle}{\|x-x^{*}\|},$
where $v\in\operatorname*{argmax}_{u\in P}\langle\nabla f(x),x-u\rangle$.
When $f$ is not convex, assuming that $f$ is weakly convex and local quadratic growth, we have the scaling inequality for a nonconvex objective function.
Proposition 2.17 (Scaling inequality for inner optima from weak convexity).
Let $P\subset\mathbb{R}^{n}$ be a nonempty compact convex set. Let $f:\mathbb{R}^{n}\to(-\infty,+\infty]$ be $\mathcal{C}^{1}$, $\rho$-weakly convex, local $\mu$-quadratic growth with $\zeta>0$ on $P$. If there exists $r>0$ so that $B(x^{*},r)\subset P$ for a minimizer $x^{*}$ of $f$ and $\rho\leq\mu$, then for all $x\in[f\leq f^{*}+\zeta]\cap P$, we have
$\langle\nabla f(x),x-v\rangle\geq r\|\nabla f(x)\|\geq\frac{r\langle\nabla f(x),x-x^{*}\rangle}{\|x-x^{*}\|},$
where $v\in\operatorname*{argmax}_{u\in P}\langle\nabla f(x),x-u\rangle$.
Proof.
We consider $x^{*}-rz$, where $z$ is a point with $\|z\|=1$ and $\langle\nabla f(x),z\rangle=\|\nabla f(x)\|$. It holds that
$\langle\nabla f(x),v\rangle\leq\langle\nabla f(x),x^{*}-rz\rangle=\langle\nabla f(x),x^{*}\rangle-r\|\nabla f(x)\|.$ (2.9)
From weak convexity, for all $x\in[f\leq f^{*}+\zeta]\cap P$, we have
$f^{*}-f(x)$ $\geq\langle\nabla f(x),x^{*}-x\rangle-\frac{\rho}{2}\|x-x^{*}\|^{2}$
$\geq\langle\nabla f(x),x^{*}-x\rangle-\frac{\rho}{\mu}(f(x)-f^{*}),$
where the last inequality holds due to the local quadratic growth condition. Because of $\rho\leq\mu$, this inequality implies
$\langle\nabla f(x),x-x^{*}\rangle\geq\left(1-\frac{\rho}{\mu}\right)(f(x)-f^{*})\geq 0.$
By rearranging (2.9) and using $\langle\nabla f(x),x-x^{*}\rangle\geq 0$, we obtain
$\langle\nabla f(x),x-v\rangle\geq\langle\nabla f(x),x-x^{*}\rangle+r\|\nabla f(x)\|\geq r\|\nabla f(x)\|.$
In addition, it holds that
$\frac{r\langle\nabla f(x),x-x^{*}\rangle}{\|x-x^{*}\|}\leq\frac{r\|\nabla f(x)\|\|x-x^{*}\|}{\|x-x^{*}\|}=r\|\nabla f(x)\|,$
where the first inequality holds because of the Cauchy–Schwarz inequality. ∎
The authors in [Lacoste-Julien2015-gb] defined a geometric distance-like constant of a polytope, known as the pyramidal width, to analyze the convergence of the away-step Frank–Wolfe algorithm (and other variants that use the pyramidal width) over polytopes. It can be interpreted as the minimal $\delta>0$ satisfying the following scaling inequality (2.10), which plays a central role in establishing convergence rates for the away-step FW algorithm; see [Braun2022-ew] for an in-depth discussion.
Lemma 2.18 (Scaling inequality via pyramidal width [Braun2022-ew, Theorem 2.26][Lacoste-Julien2015-gb, Theorem 3]).
Let $P\subset\mathbb{R}^{n}$ be a polytope and let $\delta$ denote the pyramidal width of $P$. Let $x\in P$, and let $\mathcal{S}$ denote any set of vertices of $P$ with $x\in\operatorname{conv}\mathcal{S}$. Let $\psi$ be any vector, so that we define $v^{\textup{FW}}=\operatorname*{argmin}_{v\in P}\langle\psi,v\rangle$ and $v^{\textup{A}}=\operatorname*{argmax}_{v\in\mathcal{S}}\langle\psi,v\rangle$. Then for any $y\in P$
$\langle\psi,v^{\textup{A}}-v^{\textup{FW}}\rangle\geq\delta\frac{\langle\psi,x-y\rangle}{\|x-y\|}.$ (2.10)
Note that Lemma 2.18 does not require the convexity of $f$, and we will use it for nonconvex optimization. When $\phi$ is $\sigma$-strongly convex and $\langle\psi,x-y\rangle\geq 0$, Lemma 2.18 implies
$\langle\psi,v^{\textup{A}}-v^{\textup{FW}}\rangle^{2}\geq\delta^{2}\frac{\langle\psi,x-y\rangle^{2}}{\|x-y\|^{2}}\geq\delta^{2}\sigma\frac{\langle\psi,x-y\rangle^{2}}{2D_{\phi}(x,y)}.$
Note that the last inequality of the above also holds for $\delta^{2}\sigma\frac{\langle\psi,x-y\rangle^{2}}{2D_{\phi}(y,x)}$, i.e., with $x$ and $y$ swapped in the divergence.
Geometric Hölder Error Bound Condition
We will now introduce the more compact notion of geometric Hölder error bound condition, which simply combines the pyramidal width and the Hölder error bound condition of the function $f$; see also [Braun2022-ew, Lemma 2.27] for details when $q=2$.
Lemma 2.19 (Geometric Hölder Error Bound).
Let $P$ be a polytope with pyramidal width $\delta>0$ and let $f$ be a convex function and satisfy the Hölder error bound condition with $v^{\textup{FW}}=\operatorname*{argmin}_{v\in P}\langle\nabla f(x),v\rangle$ and $v^{\textup{A}}=\operatorname*{argmax}_{v\in\mathcal{S}}\langle\nabla f(x),v\rangle$ with $\mathcal{S}\subseteq\operatorname{Vert}{P}$, so that $x\in\operatorname{conv}{\mathcal{S}}$, we have
$f(x)-f^{*}\leq\left(\frac{q}{\mu}\right)^{\frac{1}{q-1}}\left(\frac{\langle\nabla f(x),v^{\textup{A}}-v^{\textup{FW}}\rangle}{\delta}\right)^{\frac{q}{q-1}}.$
Proof.
Combining (2.10) with Lemma 2.12 for $\psi=\nabla f(x)$, we have
$f(x)-f^{*}\leq\left(\frac{q}{\mu}\right)^{\frac{1}{q-1}}\left(\frac{\langle\nabla f(x),x-x^{*}\rangle}{\|x-x^{*}\|}\right)^{\frac{q}{q-1}}\leq\left(\frac{q}{\mu}\right)^{\frac{1}{q-1}}\left(\frac{\langle\nabla f(x),v^{\textup{A}}-v^{\textup{FW}}\rangle}{\delta}\right)^{\frac{q}{q-1}}.$
From Contractions to Convergence Rates
Besides scaling inequalities and the geometric Hölder error bound, we also utilize the following lemma for convex and nonconvex optimization, which allows us to turn a contraction into a convergence rate.
Lemma 2.20 (From contractions to convergence rates [Braun2022-ew, Lemma 2.21]).
Let $\{h_{t}\}_{t}$ be a decreasing sequence of positive numbers and $c_{0}$, $c_{1}$, $c_{2}$, $\theta_{0}$ be positive numbers with $c_{1}<1$ such that $h_{1}\leq c_{0}$ and $h_{t}-h_{t+1}\geq h_{t}\min\{c_{1},c_{2}h_{t}^{\theta_{0}}\}$ for $t\geq 1$, then
$h_{t}\leq\begin{cases}c_{0}(1-c_{1})^{t-1}&\text{if}\ 1\leq t\leq t_{0},\\ \frac{(c_{1}/c_{2})^{1/\theta_{0}}}{(1+c_{1}\theta_{0}(t-t_{0}))^{1/\theta_{0}}}=\mathcal{O}(1/t^{1/\theta_{0}})&\text{if}\ t\geq t_{0},\end{cases}$
where
$t_{0}:=\max\left\{\left\lfloor\log_{1-c_{1}}\left(\frac{(c_{1}/c_{2})^{1/\theta_{0}}}{c_{0}}\right)\right\rfloor+2,1\right\}.$
In particular, we have $h_{t}\leq\epsilon$ if $t\geq t_{0}+\frac{1}{\theta_{0}c_{2}\epsilon^{\theta_{0}}}-\frac{1}{\theta_{0}c_{1}}$ and $\epsilon\leq(c_{1}/c_{2})^{1/\theta_{0}}$.
3 Proposed Algorithms
Throughout this paper, we make the following assumptions.
Assumption 3.1.
  1. (i)
    $\phi\in\mathcal{G}(C)$ with $\operatorname{cl}{C}=\operatorname{cl}{\operatorname{dom}\phi}$ is strictly convex on $C\equiv\operatorname{int}\operatorname{dom}\phi$.
  2. (ii)
    $f:\mathbb{R}^{n}\to(-\infty,+\infty]$ is a proper and lower semicontinuous function with $\operatorname{dom}\phi\subset\operatorname{dom}f$, which is $\mathcal{C}^{1}$ on $C$.
  3. (iii)
    The pair $(f,\phi)$ is $L$-smad on $P$.
  4. (iv)
    $P\subset\mathbb{R}^{n}$ is a nonempty compact convex set with $P\subset C$.
Assumption 3.1(i)-(iii) are standard for Bregman-type algorithms [Bolte2018-zt, Takahashi2024-wz]. The strict convexity ensures the existence of $\nu$ by Lemma 2.5 and is satisfied by many kernel generating distances (see Example 2.3). Assumption 3.1(iv) is also standard for FW algorithms, and the condition $P\subset C$ is natural ensuring that the Bregman distance $D_{\phi}$ is well-defined on $P$.
In what follows, let $x^{*}\in\operatorname*{argmin}_{x\in P}f(x)$ and let $f^{*}=f(x^{*})$. We recall a key property for analyzing the convergence rates of FW algorithms in convex optimization.
Lemma 3.2 (Primal gap, dual gap, and Frank–Wolfe gap [Pokutta2024-qu, Lemma 4.1]).
Let $x^{*}\in P$ be an optimal solution of (1.1) and let $f^{*}=f(x^{*})$. Suppose that Assumptions 3.1(ii), (iv) and that $f$ is convex. Then for all $x\in P$, it holds that
$f(x)-f^{*}\leq\langle\nabla f(x),x-x^{*}\rangle\leq\max_{v\in P}\langle\nabla f(x),x-v\rangle.$ (3.1)
3.1 Adaptive Bregman Step-Size Strategy
Assume that $(f,\phi)$ is $L$-smad. Substituting $x_{t+1}=(1-\gamma_{t})x_{t}+\gamma_{t}v_{t}$ with $x_{t},v_{t}\in P$ for $x_{+}$ in Lemma 2.8 provides
$f(x_{t})-f(x_{t+1})\geq\gamma_{t}\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle-L\gamma_{t}^{1+\nu}D_{\phi}(v_{t},x_{t}).$ (3.2)
Using the optimality condition of the right-hand side of (3.2) in terms of $\gamma_{t}$, we find
$\gamma_{t}=\left(\frac{\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle}{L(1+\nu)D_{\phi}(v_{t},x_{t})}\right)^{\frac{1}{\nu}}\quad\text{and}\quad f(x_{t})-f(x_{t+1})\geq\frac{\nu}{1+\nu}\frac{\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle^{1+1/\nu}}{(L(1+\nu)D_{\phi}(v_{t},x_{t}))^{1/\nu}}.$
Theoretically, when $\gamma_{t}\in[0,1]$, the Frank–Wolfe step $x_{t+1}=(1-\gamma_{t})x_{t}+\gamma_{t}v_{t}\in P$ is well-defined. We update $\gamma_{t}=\min\left\{\left(\frac{\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle}{L(1+\nu)D_{\phi}(v_{t},x_{t})}\right)^{\frac{1}{\nu}},\gamma_{\max}\right\}$ with some $\gamma_{\max}\in\mathbb{R}$ (usually set $\gamma_{\max}=1$). This step-size strategy provides the progress lemma. Note that the following lemma does not require the convexity of $f$.
Lemma 3.3 (Progress lemma from the Bregman short step-size).
Suppose that Assumption 3.1 holds. Define $x_{t+1}=(1-\gamma_{t})x_{t}+\gamma_{t}v_{t}$. Then if $x_{t+1}\in\operatorname{dom}f$,
$f(x_{t})-f(x_{t+1})\geq\frac{\nu}{1+\nu}\gamma_{t}\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle\quad\text{for}\quad 0\leq\gamma_{t}\leq\left(\frac{\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle}{L(1+\nu)D_{\phi}(v_{t},x_{t})}\right)^{\frac{1}{\nu}}.$
Proof.
Recalling (3.2), we have
$f(x_{t})-f(x_{t+1})$ $\geq\gamma_{t}\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle-L\gamma_{t}^{1+\nu}D_{\phi}(v_{t},x_{t})$
$\geq\gamma_{t}\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle-\frac{\gamma_{t}}{1+\nu}\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle$
$=\frac{\nu}{1+\nu}\gamma_{t}\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle,$
where the second inequality holds because of an upper bound of $\gamma_{t}$. ∎
If $f$ is convex, using Lemma 3.3, we have
$f(x_{t})-f(x_{t+1})\geq\frac{\nu}{1+\nu}\gamma_{t}\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle\geq\frac{\nu}{1+\nu}\gamma_{t}(f(x_{t})-f^{*}),$
which implies that the primal progress is at least $\frac{\nu}{1+\nu}$ of the Frank–Wolfe gap and of the primal gap multiplied by $\gamma_{t}$. If $\phi=\frac{1}{2}\|\cdot\|^{2}$, we have $\nu=1$ and
$\gamma_{t}=\frac{\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle}{L\|v_{t}-x_{t}\|^{2}},$
which is often called the (Euclidean) short step-size. Although the short step-size strategy does not require line searches, it requires knowledge of the value of $L$.
The exact value or the tight upper bound of $L$ is often unknown, however, the algorithm’s performance heavily depends on it; an underestimation of $L$ might lead to non-convergence, and an overestimation of $L$ might lead to slow convergence. Moreover, a worst-case $L$ might be too conversative for regimes where the function is better behaved. Due to all these reasons, Pedregosa et al. [Pedregosa2020-ay] proposed an adaptive step-size strategy for FW algorithms in the case where $\phi=\frac{1}{2}\|\cdot\|^{2}$. Vyguzov and Stonyakin [Vyguzov2024-tp] proposed a variant using Bregman distances. We have concurrently and independently formulated an improved version similar (but different in details) to [Vyguzov2024-tp], which includes updates to $\nu$ and uses different coefficients for $\gamma_{t}$. We present our algorithm in Algorithm 2.
Output : Updated estimates $\tilde{L}^{*}$ and $\tilde{\nu}^{*}$, step-size $\gamma$
[…]
2      Choose $\beta<1$, $\eta\leq 1$, and $\tau>1$
3     $\kappa\leftarrow 1$
4     $M\leftarrow\eta\tilde{L}$
5     loop
6           $\gamma\leftarrow\min\left\{\left(\frac{\langle\nabla f(x),x-v\rangle}{M(1+\kappa)D_{\phi}(v,x)}\right)^{\frac{1}{\kappa}},\gamma_{\max}\right\}$
7          if $f((1-\gamma)x+\gamma v)-f(x)-\gamma\langle\nabla f(x),v-x\rangle\leq M\gamma^{1+\kappa}D_{\phi}(v,x)$ then
8                $\tilde{L}^{*}\leftarrow M$
9               $\tilde{\nu}^{*}\leftarrow\kappa$
10               return $\tilde{L}^{*}$, $\tilde{\nu}^{*}$, $\gamma$
11                $M\leftarrow\tau M$
12               $\kappa\leftarrow\beta\kappa$
13
Algorithm 2 Adaptive Bregman step-size strategy
This step-size strategy can be used as a drop-in replacement. For example, inserting $L_{t},\nu_{t},\gamma_{t}\leftarrow\textnormal{{step\_size}}(f,\phi,x_{t},v_{t},L_{t-1},1)$ between lines 1 and 1 in Algorithm 1, we obtain that Algorithm 2 searches $L$ and $\nu$ until
$f((1-\gamma)x+\gamma v)-f(x)-\gamma\langle\nabla f(x),v-x\rangle\leq M\gamma^{1+\kappa}D_{\phi}(v,x)$ (3.3)
holds for the approximation values $M$ and $\kappa$ of $L$ and $\nu$, respectively.
Remark 3.4 (Well-definedness and termination of Algorithm 2).
By the $L$-smad property of $(f,\phi)$ and Lemma 2.8, we know that (3.3) holds for all $M\geq L$ and $\nu\geq\kappa>0$. Therefore, the loop in line 2 in Algorithm 2 is well-defined and guaranteed to terminate. Thus it suffices to update $\kappa$ in line 2 only if $D_{\phi}((1-\gamma)x+\gamma v,x)>\gamma^{1+\kappa}D_{\phi}(v,x)$ holds. Thus, $\kappa$ cannot become too small.
3.2 The Away-Step Frank–Wolfe Algorithm
In the case where $P$ is a polytope, the classical FW algorithm might zigzag when approaching the optimal face and, in consequence, converge slowly. To overcome this drawback, Wolfe [Wolfe1970-ax] introduced the away-step Frank–Wolfe algorithm and away steps, which allow the algorithm to move away from vertices in a convex combination of $x_{t}$, which can effectively shortcircuit the zigzagging. The convergence properties of this algorithm were unknown for a long time, and it was only quite recently that Lacoste-Julien and Jaggi [Lacoste-Julien2015-gb] established the linear convergence of the away-step Frank-Wolfe algorithm. Inspired by [GueLat1986-bg, Lacoste-Julien2015-gb, Wolfe1970-ax], we propose a variant of the away-step Frank–Wolfe algorithm utilizing Bregman distances as given in Algorithm 3.
For the following discussion, we introduce the following notions. In the same way [Braun2022-ew], we define an active set $\mathcal{S}\subset\operatorname{Vert}P$ and the away vertex as $v_{t}^{\textup{A}}\in\operatorname*{argmax}_{v\in\mathcal{S}}\langle\nabla f(x_{t}),v\rangle$, where $x_{t}$ is a (strict) convex combination of elements in $\mathcal{S}$, i.e., $x_{t}=\sum_{v\in\mathcal{S}}\lambda_{v}x_{t}$ with $\lambda_{v}>0$ for all $v\in\mathcal{S}$ and $\sum_{v\in\mathcal{S}}\lambda_{v}=1$. Algorithm 3 updates $\nu$ until $D_{\phi}((1-\gamma)x_{t}+\gamma v_{t},x_{t})\leq\gamma^{1+\nu}D_{\phi}(v,x)$ holds in lines 3-3. If $\nu$ is known, one only updates $\gamma_{t}$. If $\nu$ and $L$ are unknown, we can use $L_{t},\nu_{t},\gamma_{t}\leftarrow\textnormal{{step\_size}}(f,\phi,x_{t},v_{t},L_{t-1},\gamma_{t,\max})$ with $\gamma\leftarrow\min\left\{\left(\frac{\langle\nabla f(x_{t}),d_{t}\rangle}{M(1+\kappa)D_{\phi}(v_{t},x_{t})}\right)^{\frac{1}{\kappa}},\gamma_{t,\max}\right\}$ instead of line 2 in Algorithm 2. When $\phi=\frac{1}{2}\|\cdot\|^{2}$ and $\nu=1$, Algorithm 3 corresponds to the classical away-step FW algorithm.
Input: Initial point $x_{0}\in\operatorname*{argmin}_{v\in P}\langle\nabla f(x),v\rangle$ for $x\in P$, $\beta<1$
1 $\mathcal{S}_{0}\leftarrow\{x_{0}\}$, $\lambda_{x_{0},0}\leftarrow 1$
2for $t=0,\ldots$ do
3      $v_{t}^{\textup{FW}}\leftarrow\operatorname*{argmin}_{v\in P}\langle\nabla f(x_{t}),v\rangle$
4     $v_{t}^{\textup{A}}\leftarrow\operatorname*{argmax}_{v\in\mathcal{S}_{t}}\langle\nabla f(x_{t}),v\rangle$
5     if $\langle\nabla f(x_{t}),x_{t}-v_{t}^{\textup{FW}}\rangle\geq\langle\nabla f(x_{t}),v_{t}^{\textup{A}}-x_{t}\rangle$ then
6           $v_{t}\leftarrow v_{t}^{\textup{FW}}$, $d_{t}\leftarrow x_{t}-v_{t}^{\textup{FW}}$, $\gamma_{t,\max}\leftarrow 1$
7           else
8                $v_{t}\leftarrow v_{t}^{\textup{A}}$, $d_{t}\leftarrow v_{t}^{\textup{A}}-x_{t}$, $\gamma_{t,\max}\leftarrow\frac{\lambda_{v_{t}^{\textup{A}},t}}{1-\lambda_{v_{t}^{\textup{A}},t}}$
9                $\nu\leftarrow 1$
10               while $D_{\phi}((1-\gamma)x_{t}+\gamma v_{t},x_{t})\leq\gamma^{1+\nu}D_{\phi}(v_{t},x_{t})$ do
11                     $\gamma\leftarrow\min\left\{\left(\frac{\langle\nabla f(x_{t}),d_{t}\rangle}{L(1+\nu)D_{\phi}(v_{t},x_{t})}\right)^{\frac{1}{\nu}},\gamma_{t,\max}\right\}$, $\nu\leftarrow\beta\nu$
12
13                    $\gamma_{t}\leftarrow\gamma$, $x_{t+1}\leftarrow x_{t}-\gamma_{t}d_{t}$
14                    if $\langle\nabla f(x_{t}),x_{t}-v_{t}^{\textup{FW}}\rangle\geq\langle\nabla f(x_{t}),v_{t}^{\textup{A}}-x_{t}\rangle$ then
15                          $\lambda_{v,t+1}\leftarrow(1-\gamma_{t})\lambda_{v,t}$ for all $v_{t}\in\mathcal{S}_{t}\setminus\{v_{t}^{\textup{FW}}\}$
16                         $\lambda_{v_{t}^{\textup{FW}},t+1}\leftarrow\begin{cases}\gamma_{t}&\text{if}\ v_{t}^{\textup{FW}}\notin\mathcal{S}_{t}\\ (1-\gamma_{t})\lambda_{v_{t}^{\textup{FW}},t}+\gamma_{t}&\text{if}\ v_{t}^{\textup{FW}}\in\mathcal{S}_{t}\end{cases}$
17                         $\mathcal{S}_{t+1}\leftarrow\begin{cases}\mathcal{S}_{t}\cup\{v_{t}^{\textup{FW}}\}&\text{if}\ \gamma_{t}<1\\ \{v_{t}^{\textup{FW}}\}&\text{if}\ \gamma_{t}=1\end{cases}$
18                          else
19                               $\lambda_{v,t+1}\leftarrow(1+\gamma_{t})\lambda_{v,t}$ for all $v_{t}\in\mathcal{S}_{t}\setminus\{v_{t}^{\textup{A}}\}$
20                              $\lambda_{v_{t}^{\textup{A}},t+1}\leftarrow(1+\gamma_{t})\lambda_{v_{t}^{\textup{A}},t}-\gamma_{t}$
21                              $\mathcal{S}_{t+1}\leftarrow\begin{cases}\mathcal{S}_{t}\setminus\{v_{t}^{\textup{A}}\}&\text{if}\ \lambda_{v_{t}^{\textup{A}},t+1}=0\\ \mathcal{S}_{t}&\text{if}\ \lambda_{v_{t}^{\textup{A}},t+1}>0\end{cases}$
22
Algorithm 3 Away-Step Frank–Wolfe algorithm with the Bregman distance
4 Convergence Analysis for Convex Optimization
In this section, we assume that $f$ is convex.
Assumption 4.1.
The objective function $f$ is convex.
4.1 Sublinear Convergence
We establish a sublinear convergence rate of the FW algorithm under the smooth adaptable property. It is a similar result to [Vyguzov2024-tp, Theorem 1]. We conducted its proof following [Braun2022-ew, Jaggi2013-oz, Pokutta2024-qu]. The FW algorithm uses the open loop step-size, i.e., $\gamma_{t}=\frac{2}{2+t}$ in the following theorem.
Theorem 4.2 (Primal convergence of the Frank–Wolfe algorithm).
Suppose that Assumptions 3.1 and 4.1 hold. Let $D:=\sqrt{\sup_{x,y\in P}D_{\phi}(x,y)}$ be the diameter of $P$ characterized by the Bregman distance $D_{\phi}$. Consider the iterates of Algorithm 1 with the open loop step-size, i.e., $\gamma_{t}=\frac{2}{2+t}$. Then, it holds that, for all $t\geq 1$,
$f(x_{t})-f^{*}\leq\frac{2^{1+\nu}LD^{2}}{(t+2)^{\nu}},$ (4.1)
and hence for any accuracy $\epsilon>0$ we have $f(x_{t})-f^{*}\leq\epsilon$ for all $t\geq\left(\frac{2^{1+\nu}LD^{2}}{\epsilon}\right)^{\frac{1}{\nu}}$.
Proof.
For some $\nu\in(0,1]$, we have
$f(x_{t})-f(x_{t+1})$ $\geq\gamma_{t}\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle-L\gamma_{t}^{1+\nu}D_{\phi}(v_{t},x_{t})$
$\geq\gamma_{t}(f(x_{t})-f^{*})-L\gamma_{t}^{1+\nu}D_{\phi}(v_{t},x_{t}),$
where the first inequality holds because of Lemma 2.8 and the last inequality holds because of Lemma 3.2. Subtracting $f^{*}$ on both sides, using $D_{\phi}(v_{t},x_{t})\leq D^{2}$, and rearranging leads to
$f(x_{t+1})-f^{*}\leq(1-\gamma_{t})(f(x_{t})-f^{*})+L\gamma_{t}^{1+\nu}D^{2}.$
When $t=0$, it follows $f(x_{1})-f^{*}\leq LD^{2}\leq 2LD^{2}$. Now, we consider $t\geq 1$ and obtain
$f(x_{t+1})-f^{*}$ $\leq(1-\gamma_{t})(f(x_{t})-f^{*})+L\gamma_{t}^{1+\nu}D^{2}$
$\leq\frac{t}{2+t}(f(x_{t})-f^{*})+\frac{2^{1+\nu}}{(2+t)^{1+\nu}}LD^{2}$
$\leq\frac{t}{2+t}\frac{2^{1+\nu}LD^{2}}{(2+t)^{\nu}}+\frac{2^{1+\nu}}{(2+t)^{1+\nu}}LD^{2}$
$=\frac{2^{1+\nu}LD^{2}}{(3+t)^{\nu}}\left(\frac{(3+t)^{\nu}(1+t)}{(2+t)^{1+\nu}}\right)\leq\frac{2^{1+\nu}LD^{2}}{(3+t)^{\nu}},$
where the last inequality holds due to $(3+t)^{\nu}(1+t)\leq(2+t)^{1+\nu}$ with $0<\nu\leq 1$. ∎
If $\phi=\frac{1}{2}\|\cdot\|^{2}$ and $\nu=1$ in (4.1), we have
$f(x_{t})-f^{*}\leq\frac{4LD^{2}}{t+2},$
which is the same as a sublinear convergence rate of the classical FW algorithm.
Remark 4.3.
While the convergence rate (4.1) is the same as [Vyguzov2024-tp], Vyguzov and Stonyakin assume that the triangle scaling property holds for $D_{\phi}$. In contrast, we require significantly weaker assumptions: it is enough to assume that $\phi$ is strictly convex due to Lemma 2.5 in order to establish Theorem 4.2.
4.2 Accelerated Convergence
We will now establish accelerated convergence rates better than $O(1/t)$ up to linear convergence depending on the choice of parameters. First, we establish the accelerated convergence of Algorithm 1 with the Bregman short step-size, i.e.,
$\gamma_{t}=\min\left\{\left(\frac{\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle}{L(1+\nu)D_{\phi}(v_{t},x_{t})}\right)^{\frac{1}{\nu}},1\right\},$
in the case where the optimal solution lies in the relative interior.
Theorem 4.4 (Accelerated convergence for inner optima).
Suppose that Assumptions 3.1 and 4.1 hold. Let $f$ satisfy the Hölder error bound condition with $q\geq 1+\nu$ and $\mu>0$ and $D:=\sqrt{\sup_{x,y\in P}D_{\phi}(x,y)}$ be the diameter of $P$. Assume that there exists a minimizer $x^{*}\in\operatorname{int}P$, i.e., there exists an $r>0$ with $B(x^{*},r)\subset P$. Consider the iterates of Algorithm 1 with $\gamma_{t}=\min\left\{\left(\frac{\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle}{L(1+\nu)D_{\phi}(v_{t},x_{t})}\right)^{\frac{1}{\nu}},1\right\}$. Then, it holds that
$f(x_{t})-f^{*}\leq\begin{cases}h_{t}\leq\max\left\{\frac{1}{1+\nu},1-\frac{\nu}{1+\nu}\frac{r^{1+1/\nu}}{c^{1+1/\nu}D^{2/\nu}}\right\}^{t-1}LD^{2}&\text{if}\ q=1+\nu,\\ \frac{LD^{2}}{(1+\nu)^{t-1}}&\text{if}\ 1\leq t\leq t_{0},q>1+\nu,\\ \frac{(L(1+\nu)(c/r)^{1+\nu}D^{2})^{q/(q-1-\nu)}}{\left(1+\frac{1-\nu}{2(1+\nu)}(t-t_{0})\right)^{\nu q/(q-1-\nu)}}=\mathcal{O}(1/t^{\nu q/(q-1-\nu)})&\text{if}\ t\geq t_{0},q>1+\nu,\end{cases}$
for all $t\geq 1$ where $c=(q/\mu)^{1/q}$ and
$t_{0}:=\max\left\{\left\lfloor\log_{\frac{1}{1+\nu}}\left(\frac{(L(1+\nu(c/r)^{1+\nu}D^{2})^{q/(q-1-\nu)}}{LD^{2}}\right)\right\rfloor+2,1\right\}.$
Proof.
Let $h_{t}:=f(x_{t})-f^{*}$. Using Lemma 3.3, we have
$h_{t}-h_{t+1}=f(x_{t})-f(x_{t+1})\geq\frac{\nu}{1+\nu}\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle\gamma_{t},$
where $\gamma_{t}=\min\left\{\left(\frac{\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle}{L(1+\nu)D_{\phi}(v_{t},x_{t})}\right)^{\frac{1}{\nu}},1\right\}$. We consider two cases: (i) $\gamma_{t}<1$ and (ii) $\gamma_{t}=1$.
(i) $\gamma_{t}<1$: Using $\gamma_{t}=\left(\frac{\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle}{L(1+\nu)D_{\phi}(v_{t},x_{t})}\right)^{\frac{1}{\nu}}$, we have
$h_{t}-h_{t+1}$ $\geq\frac{\nu}{1+\nu}\frac{\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle^{1+1/\nu}}{(L(1+\nu)D_{\phi}(v_{t},x_{t}))^{1/\nu}}$
$\geq\frac{\nu}{1+\nu}\frac{r^{1+1/\nu}\|\nabla f(x_{t})\|^{1+1/\nu}}{(L(1+\nu))^{1/\nu}D^{2/\nu}}$
$\geq\frac{\nu}{1+\nu}\frac{r^{1+1/\nu}}{Mc^{1+1/\nu}D^{2/\nu}}h_{t}^{\frac{(1+\nu)(q-1)}{\nu q}},$
where the second inequality holds from $\langle f(x_{t}),x_{t}-v_{t}\rangle\geq r\|\nabla f(x_{t})\|$ in Proposition 2.16, and the last inequality holds because $f$ satisfies the $c$-Łojasiewicz inequality with $c=(q/\mu)^{1/q}$ (see Remark 2.13), and $M:=(L(1+\nu))^{1/\nu}$.
(ii) $\gamma_{t}=1$: Using (3.1), we have
$h_{t}-h_{t+1}=f(x_{t})-f(x_{t+1})\geq\frac{\nu}{1+\nu}\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle\geq\frac{\nu}{1+\nu}(f(x_{t})-f^{*})=\frac{\nu}{1+\nu}h_{t},$
where the second inequality holds due to the convexity of $f$.
From (i) and (ii), we have
$h_{t}-h_{t+1}\geq\min\left\{\frac{\nu}{1+\nu}h_{t},\frac{\nu}{1+\nu}\frac{r^{1+1/\nu}}{Mc^{1+1/\nu}D^{2/\nu}}h_{t}^{\frac{(1+\nu)(q-1)}{\nu q}}\right\}.$
When $q=1+\nu$, we have
$h_{t}-h_{t+1}\geq\min\left\{\frac{\nu}{1+\nu},\frac{\nu}{1+\nu}\frac{r^{1+1/\nu}}{Mc^{1+1/\nu}D^{2/\nu}}\right\}\cdot h_{t}.$
This inequality and the initial bound $f(x_{1})-f^{*}\leq LD^{2}$ due to Lemma 2.7 imply
$h_{t}\leq\max\left\{\frac{1}{1+\nu},1-\frac{\nu}{1+\nu}\frac{r^{1+1/\nu}}{Mc^{1+1/\nu}D^{2/\nu}}\right\}^{t-1}LD^{2}.$
On the other hand, when $q>1+\nu$, we have
$h_{t}-h_{t+1}\geq\min\left\{\frac{\nu}{1+\nu},\frac{\nu}{1+\nu}\frac{r^{1+1/\nu}}{Mc^{1+1/\nu}D^{2/\nu}}h_{t}^{\frac{q-1-\nu}{\nu q}}\right\}\cdot h_{t}.$
Using $f(x_{1})-f^{*}\leq LD^{2}$ and Lemma 2.20 with $c_{0}=LD^{2}$, $c_{1}=\frac{\nu}{1+\nu}$, $c_{2}=\frac{\nu}{1+\nu}\frac{r^{1+1/\nu}}{Mc^{1+1/\nu}D^{2/\nu}}$, and $\theta_{0}=\frac{q-1-\nu}{\nu q}>0$ from $q>1+\nu$, we have the claim. ∎
Because Algorithm 2 is well-defined (see Remark 3.4), the convergence result of the FW algorithm using the adaptive step-size strategy (Algorithm 2) is essentially the same as the one from Theorem 4.4 which uses Bregman short steps; up to small errors arising from the approximation whose precise analysis we skip for the sake of brevity. Note that Vyguzov and Stonyakin [Vyguzov2024-tp, Theorem 2] also established linear convergence, assuming that $f$ is relatively strongly convex, the optimal solution belongs to int $P$, the angle condition holds, and the triangle scaling property holds. However, these requirements are stronger than ours here, and they do not establish sublinear but accelerated rates for intermediate parameter regimes.
Next, we establish the linear convergence of Algorithm 3. Recall that the pyramidal width is the minimal $\delta>0$ satisfying Lemma 2.18 (see [Braun2022-ew, Lemma 2.26] or [Lacoste-Julien2015-gb, Theorem 3] for an in-depth discussion). We make the following assumption:
Assumption 4.5.
The kernel generating distance $\phi$ is $\sigma$-strongly convex.
Under Assumption 4.5, it holds that $\delta\leq D$, and this will be important in establishing convergence rates of Algorithm 3.
Theorem 4.6 (Accelerated convergence of the away-step FW algorithm).
Suppose that Assumptions 3.14.1, and 4.5 hold. Let $P\subset\mathbb{R}^{n}$ be a polytope and $f$ satisfy the Hölder error bound condition with $q>1+\nu$ or $(\nu,q)=(1,2)$. The convergence rate of Algorithm 3 is linear: for all $t\geq 1$
$f(x_{t})-f^{*}\leq\begin{cases}\left(1-\frac{\mu}{32L}\frac{\delta^{2}}{D^{2}}\right)^{\lceil(t-1)/2\rceil}LD^{2}&\text{if}\ (\nu,q)=(1,2),\\ \frac{1}{(1+\nu)^{\lceil(t-1)/2\rceil}}LD^{2}&\text{if}\ 1\leq t\leq t_{0},q>1+\nu,\\ \frac{(L(1+\nu)D^{2}/((\mu/q)^{1/q}\delta/2)^{1+\nu})^{q/(q-1-\nu)}}{\left(1+\frac{1-\nu}{2(1+\nu)}\lceil(t-t_{0})/2\rceil\right)^{\nu q/(q-1-\nu)}}=\mathcal{O}(1/t^{\nu q/(q-1-\nu)})&\text{if}\ t\geq t_{0},q>1+\nu,\end{cases}$
where $D:=\sqrt{\sup_{x,y\in P}D_{\phi}(x,y)}$ and $\delta$ are the diameter and the pyramidal width of the polytope $P$, respectively, and
$t_{0}:=\max\left\{\left\lfloor\log_{\frac{1}{1+\nu}}\frac{(L(1+\nu)D^{2}/((\mu/q)^{1/q}\delta/2)^{1+\nu})^{q/(q-1-\nu)}}{LD^{2}}\right\rfloor+2,1\right\}.$
Proof.
By using the induced guarantee on the primal gap via Lemma 2.19, we have
$h_{t}=f(x_{t})-f^{*}\leq\left(\frac{q}{\mu}\right)^{\frac{1}{q-1}}\left(\frac{\langle\nabla f(x_{t}),v_{t}^{\textup{A}}-v_{t}^{\textup{FW}}\rangle}{\delta}\right)^{\frac{q}{q-1}}\leq\left(\frac{q}{\mu}\right)^{\frac{1}{q-1}}\left(\frac{2\langle\nabla f(x_{t}),d_{t}\rangle}{\delta}\right)^{\frac{q}{q-1}},$ (4.2)
where the last inequality holds because $d_{t}$ is either $x_{t}-v_{t}^{\textup{FW}}$ or $v_{t}^{\textup{A}}-x_{t}$ with $\langle\nabla f(x_{t}),d_{t}\rangle\geq\langle\nabla f(x),v^{\textup{A}}-v^{\textup{FW}}\rangle/2$ in Lines 3 and 3 of Algorithm 3. We obtain
$h_{t}-h_{t+1}$ $\geq\frac{\nu}{1+\nu}\langle\nabla f(x_{t}),d_{t}\rangle\min\left\{\gamma_{t,\max},\left(\frac{\langle\nabla f(x_{t}),d_{t}\rangle}{L(1+\nu)D_{\phi}(v_{t},x_{t})}\right)^{\frac{1}{\nu}}\right\}$
$=\min\left\{\gamma_{t,\max}\frac{\nu}{1+\nu}\langle\nabla f(x_{t}),d_{t}\rangle,\frac{\nu}{1+\nu}\frac{\langle\nabla f(x_{t}),d_{t}\rangle^{1+1/\nu}}{(L(1+\nu)D_{\phi}(v_{t},x_{t}))^{1/\nu}}\right\}$
$\geq\min\left\{\gamma_{t,\max}\frac{\nu h_{t}}{1+\nu},\frac{\nu}{1+\nu}\frac{((\mu/q)^{1/q}h_{t}^{1-1/q}\delta/2)^{\frac{1+\nu}{\nu}}}{(L(1+\nu)D^{2})^{1/\nu}}\right\}$
$=\min\left\{\frac{\nu h_{t}}{1+\nu}\gamma_{t,\max},\frac{\nu}{1+\nu}\frac{((\mu/q)^{1/q}\delta/2)^{\frac{1+\nu}{\nu}}}{(L(1+\nu)D^{2})^{1/\nu}}h_{t}^{\frac{(1+\nu)(q-1)}{\nu q}}\right\}$
where the first inequality holds due to Lemma 3.3 with $\gamma_{t,\max}=1$ (Frank–Wolfe steps) and $\gamma_{t,\max}=\frac{\lambda_{v_{t}^{\textup{A}},t}}{1-\lambda_{v_{t}^{\textup{A}},t}}$ (away steps), and the second inequality holds because of (4.2) and $\langle\nabla f(x_{t}),d_{t}\rangle\geq\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle\geq h_{t}$.
In $(\nu,q)=(1,2)$, for Frank–Wolfe steps, $\gamma_{t,\max}=1\geq\mu\delta^{2}/LD^{2}\geq\mu\delta^{2}/32LD^{2}$. For away steps, we only rely on monotone progress $h_{t+1}<h_{t}$ because it is difficult to estimate $\gamma_{t,\max}$ below. However, $\gamma_{t,\max}$ cannot be small too often, which is the key point. Let us consider $\gamma_{t}=\gamma_{t,\max}$ in an away step. In that case, $v^{\textup{A}}_{t}$ is removed from the active set $\mathcal{S}_{t+1}$ in line 3. Moreover, the active set can only grow in Frank–Wolfe steps (line 3). It is impossible to remove more vertices from $\mathcal{S}_{t+1}$ than have been added in Frank–Wolfe steps. Therefore, at most half of iterations until $t$ iterations are in away steps, i.e., in all other steps, we have $\gamma_{t}=\frac{\langle\nabla f(x_{t}),d_{t}\rangle}{2LD_{\phi}(v_{t},x_{t})}<\gamma_{t,\max}$ and then $h_{t}-h_{t+1}\geq\frac{\mu\delta^{2}}{32LD^{2}}h_{t}$. Because we have $h_{t+1}\leq\left(1-\frac{\mu\delta^{2}}{32LD^{2}}\right)h_{t}$ for at least half of the iterations and $h_{t+1}\leq h_{t}$ for the rest, we obtain
$f(x_{t})-f^{*}\leq\left(1-\frac{\mu}{32L}\frac{\delta^{2}}{D^{2}}\right)^{\lceil(t-1)/2\rceil}LD^{2}.$
In $q>1+\nu$, we have $h_{t}-h_{t+1}\geq\min\left\{\frac{\nu}{1+\nu},\frac{\nu}{1+\nu}\frac{((\mu/q)^{1/q}\delta/2)^{\frac{1+\nu}{\nu}}}{(L(1+\nu)D^{2})^{1/\nu}}h_{t}^{\frac{q-1-\nu}{\nu q}}\right\}\cdot h_{t}$ for at least half of the iterations (Frank–Wolfe steps) and $h_{t+1}\leq h_{t}$ for the rest (away steps). The initial bound $f(x_{1})-f^{*}\leq LD^{2}$ holds generally for the Frank–Wolfe algorithm (see the proof of Theorem 4.2 and [Braun2022-ew, Remark 2.4]). We use Lemma 2.20 with $c_{0}=LD^{2}$, $c_{1}=\frac{\nu}{1+\nu}$, $c_{2}=c_{1}\frac{((\mu/q)^{1/q}\delta/2)^{\frac{1+\nu}{\nu}}}{(L(1+\nu)D^{2})^{1/\nu}}$, and $\theta_{0}=\frac{q-1-\nu}{\nu q}>0$ from $q>1+\nu$ and obtain the claim. ∎
Remark 4.7 (Compatibility of parameters).
The condition $q>1+\nu$ is necessary in case that $(\nu,q)\neq(1,2)$. The reason is as follows. In the case $q=1+\nu$, for Frank–Wolfe steps, we have
$\frac{((\mu/q)^{1/q}\delta/2)^{\frac{1+\nu}{\nu}}}{(L(1+\nu)D^{2})^{1/\nu}}=\left(\frac{\mu}{L}\right)^{1/\nu}\left(\frac{\delta}{D}\right)^{\frac{1+\nu}{\nu}}\frac{1}{2^{\frac{1+\nu}{\nu}}(1+\nu)^{2/\nu}D^{(1-\nu)/\nu}},$ (4.3)
which might be greater than 1 because $\frac{1}{D^{(1-\nu)/\nu}}\geq 1$ if $D<1$ and $\nu$ are small enough. In order to make (4.3) smaller than 1, $\nu$ should be 1, i.e., $(\nu,q)=(1,2)$. When $\nu=1$ and $q\neq 2$, it reduces to the $q>1+\nu$ case.
5 Convergence Analysis for Nonconvex Optimization
In this section, we consider a nonconvex objective function. We establish global sublinear convergence to a stationary point of (1.1), i.e., $x\in P$ such that $\max_{v\in P}\langle\nabla f(x),x-v\rangle=0$. We also obtain local sublinear and local linear convergence to a minimizer of (1.1).
5.1 Global Convergence
We show that Algorithm 1 with $\gamma_{t}=\gamma:=1/(T+1)^{\frac{1}{1+\nu}}$ globally converges to a stationary point, where $T\in\mathbb{N}$ is the number of iterations. Its proof is inspired by [Lacoste-Julien2016-mj, Theorem 1] and [Pokutta2024-qu, Theorem 4.7] and identical to those for the case when $\phi=\frac{1}{2}\|\cdot\|^{2}$.
Theorem 5.1 (Global sublinear convergence for nonconvex optimization).
Suppose that Assumption 3.1 holds. Let $D:=\sqrt{\sup_{x,y\in P}D_{\phi}(x,y)}$ be the diameter of $P$ characterized by $D_{\phi}$ and let $T\in\mathbb{N}$. Then, the iterates of the FW algorithm with $\gamma_{t}=\gamma:=1/(T+1)^{\frac{1}{1+\nu}}$ satisfy
$G_{T}:=\min_{0\leq t\leq T}\max_{v_{t}\in P}\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle\leq\frac{2\max\{h_{0},LD^{2}\}}{(T+1)^{\frac{\nu}{1+\nu}}},$
where $h_{0}=f(x_{0})-f^{*}$ is the primal gap at $x_{0}$.
Proof.
Substituting $x_{t+1}$ for $x_{+}$ and $x_{t}$ for $x$ in Lemma 2.8, we have the following inequality:
$f(x_{t})-f(x_{t+1})$ $\geq\gamma\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle-L\gamma^{1+\nu}D_{\phi}(v_{t},x_{t}).$
Summing up the above inequality along $t=1,\ldots,T$ and rearranging provides
$\gamma\sum_{t=0}^{T}\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle$ $\leq f(x_{0})-f(x_{T+1})+\gamma^{1+\nu}\sum_{t=0}^{T}LD_{\phi}(x_{t},v_{t})$
$\leq f(x_{0})-f^{*}+\gamma^{1+\nu}\sum_{t=0}^{T}LD^{2}=h_{0}+\gamma^{1+\nu}(T+1)LD^{2}.$
Dividing by $\gamma(T+1)$ on the both sides, we obtain
$G_{T}\leq\frac{1}{T+1}\sum_{t=0}^{T}\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle\leq\frac{h_{0}}{\gamma(T+1)}+\gamma^{\nu}LD^{2},$
which, for $\gamma=1/(T+1)^{\frac{1}{1+\nu}}$, implies
$G_{T}\leq\frac{1}{T+1}\sum_{t=0}^{T}\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle\leq(h_{0}+LD^{2})(T+1)^{-\frac{\nu}{1+\nu}}\leq 2\max\{h_{0},LD^{2}\}(T+1)^{-\frac{\nu}{1+\nu}}.$
This is the desired claim. ∎
As mentioned above, we generalize prior similar results. In fact, in the case where $\phi=\frac{1}{2}\|\cdot\|^{2}$ and $\nu=1$, we have $\gamma_{t}=\frac{1}{\sqrt{T+1}}$ and obtain as guarantee
$\min_{0\leq t\leq T}\max_{v_{t}\in P}\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle\leq\frac{2\max\{h_{0},LD^{2}\}}{\sqrt{T+1}},$
which is the same rate as [Lacoste-Julien2016-mj, Theorem 1].
5.2 Local Convergence
Next, we show that the FW algorithm converges to a minimizer $x^{*}\in\operatorname*{argmin}_{x\in P}f(x)$ when an initial point is close enough to $x^{*}$. We need the weak convexity of $f$ for its proof.
Assumption 5.2.
  1. (i)
    $f$ is $\rho$-weakly convex.
  2. (ii)
    $f$ is local $\mu$-quadratic growth with $\zeta>0$.
Under Assumption 5.2(i), the following primal gap lemma holds, whose proof is immediate.
Lemma 5.3 (Primal gap, dual gap, and Frank–Wolfe gap for weakly convex functions).
Suppose that Assumptions 3.1(ii)(iv), and 5.2(i). Let $x^{*}\in P$ an optimal solution of (1.1) and let $f^{*}=f(x^{*})$. For all $x\in P$, it holds:
$f(x)-f^{*}$ $\leq\langle\nabla f(x),x-x^{*}\rangle+\frac{\rho}{2}\|x-x^{*}\|^{2}$
$\leq\max_{v\in P}\langle\nabla f(x),x-v\rangle+\frac{\rho}{2}\|x-x^{*}\|^{2}.$ (5.1)
Proof.
The first inequality follows from the weak convexity of $f$ and the second follows from the maximality of $\langle\nabla f(x),x-v\rangle$. ∎
Note that Lemma 5.3 differs from Lemma 3.2 in the additional term $\frac{\rho}{2}\|x-x^{*}\|^{2}$. Moreover, we prove a lemma used in the proof of convergence analysis.
Lemma 5.4.
For any $\nu\in(0,1]$ and $t\geq 0$, it holds that
$h(t):=\frac{(t+3)^{\nu}(t+2-\nu)}{(t+2)^{1+\nu}}\leq 1.$ (5.2)
Proof.
We have
$\lim_{t\to\infty}h(t)=\lim_{t\to\infty}\frac{(t+3)^{\nu}(t+2-\nu)}{(t+2)^{1+\nu}}=\lim_{t\to\infty}\left(1+\frac{1}{t+2}\right)^{\nu}\left(1-\frac{\nu}{t+2}\right)=1.$
We take the logarithm of $h(t)$, that is, $\log h(t)=\nu\log(t+3)+\log(t+2-\nu)-(1+\nu)\log(t+2)$ and obtain
$\frac{h^{\prime}(t)}{h(t)}$ $=\frac{\nu}{t+3}+\frac{1}{t+2-\nu}-\frac{1+\nu}{t+2}=\frac{-2\nu^{2}+10\nu}{(t+2)(t+3)(t+2-\nu)}>0,$
where the last inequality holds for $\nu\in(0,1]$ and $t\geq 0$. Since $h(t)>0$ for $t\geq 0$, the above inequality implies $h^{\prime}(t)>0$. Therefore, we have $\sup h(t)=1$, which implies $h(t)\leq 1$. ∎
Now we show sublinear convergence with $\gamma_{t}=\frac{2}{2+t}$. The proof is a modified version of Theorem 4.2.
Theorem 5.5 (Local sublinear convergence).
Suppose that Assumptions 3.14.5, and 5.2 hold. Let $D:=\sqrt{\sup_{x,y\in P}D_{\phi}(x,y)}$ be the diameter of $P$ characterized by the Bregman distance $D_{\phi}$. Consider the iterates of Algorithm 1 with the open loop step-size, i.e., $\gamma_{t}=\frac{2}{2+t}$. Then, if $\rho/\sigma\leq L$ and $\frac{3\rho}{\mu}\leq 2-\nu$ hold, it holds that, for all $t\geq 1$,
$f(x_{t})-f^{*}\leq\frac{2^{1+\nu}\mu LD^{2}}{\rho(t+2)^{\nu}},$ (5.3)
and hence for any accuracy $\epsilon>0$ we have $f(x_{t})-f^{*}\leq\epsilon$ for all $t\geq\left(\frac{2^{1+\nu}\mu LD^{2}}{\rho\epsilon}\right)^{\frac{1}{\nu}}$.
Proof.
Using Lemma 2.8, we have
$f(x_{t})-f(x_{t+1})$ $\geq\gamma_{t}\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle-L\gamma_{t}^{1+\nu}D_{\phi}(v_{t},x_{t})$
$\geq\gamma_{t}\left(f(x_{t})-f^{*}-\frac{\rho}{2}\|x_{t}-x^{*}\|^{2}\right)-L\gamma_{t}^{1+\nu}D_{\phi}(v_{t},x_{t}),$
where the last inequality holds because of (5.1) in Lemma 5.3. Subtracting $f^{*}$ and rearranging provides
$f(x_{t+1})-f^{*}$ $\leq(1-\gamma_{t})(f(x_{t})-f^{*})+\frac{\rho\gamma_{t}}{2}\|x_{t}-x^{*}\|^{2}+L\gamma_{t}^{1+\nu}D_{\phi}(v_{t},x_{t})$ (5.4)
$\leq\left(1-\left(1-\frac{\rho}{\mu}\right)\gamma_{t}\right)(f(x_{t})-f^{*})+L\gamma_{t}^{1+\nu}D^{2},$ (5.5)
where the last inequality holds due to the quadratic growth condition. For $t=0$, using (5.4), we have
$f(x_{1})-f^{*}\leq\frac{\rho}{2}\|x_{0}-x^{*}\|^{2}+LD_{\phi}(v_{0},x_{0})\leq\left(\frac{\rho}{\sigma}+L\right)D^{2}\leq 2LD^{2}\leq\frac{2\mu}{\rho}LD^{2},$ (5.6)
where the second inequality holds because $\phi$ is $\sigma$-strongly convex and the last inequality holds because of $3<6/(2-\nu)\leq 2\mu/\rho$. Now we consider $t\geq 1$. Using (5.5) and $\gamma_{t}=\frac{2}{2+t}$, we have
$f(x_{t+1})-f^{*}$ $\leq\left(t+\frac{2\rho}{\mu}\right)\frac{2^{1+\nu}\mu LD^{2}}{\rho(t+2)^{1+\nu}}+\frac{2^{1+\nu}LD^{2}}{(t+2)^{1+\nu}}$
$=\left(t+\frac{2\rho}{\mu}+\frac{\rho}{\mu}\right)\frac{\mu}{\rho}\frac{2^{1+\nu}LD^{2}}{(t+2)^{1+\nu}}$
$=\frac{\mu}{\rho}\frac{2^{1+\nu}LD^{2}}{(t+3)^{\nu}}\left(\frac{(t+3)^{\nu}(t+3\rho/\mu)}{(t+2)^{1+\nu}}\right)$
$\leq\frac{2^{1+\nu}\mu LD^{2}}{\rho(t+3)^{\nu}},$
where the last inequality holds because of $(t+3)^{\nu}(t+3\rho/\mu)\leq(t+3)^{\nu}(t+2-\nu)\leq(t+2)^{1+\nu}$ from (5.2). ∎
We can apply the quadratic growth condition again as before and obtain
$\operatorname{dist}(x_{t},\mathcal{X}^{*})^{2}\leq\frac{2}{\mu}(f(x_{t})-f^{*})\leq\frac{2^{2+\nu}LD^{2}}{\rho(t+2)^{\nu}},$
and hence
$\operatorname{dist}(x_{t},\mathcal{X}^{*})\leq\frac{2^{1+\nu/2}D\sqrt{L}}{\sqrt{\rho}(t+2)^{\nu/2}}.$
Note that Theorem 5.5 does not require knowledge of the number of iterations $T$ ahead of time compared to Theorem 5.1. We stress, nonetheless, that the latter can also be adjusted using a different step-size strategy to obtain an any-time guarantee; see [Braun2022-ew] for details for the standard Euclidean case, which can be generalized to our setup. Moreover, we can apply Theorem 5.5 to the classical FW algorithm, i.e., $\phi=\frac{1}{2}\|\cdot\|^{2}$ and $\nu=1$. Using (5.3), $D_{\textup{Euc}}:=\sup_{x,y\in P}\|x-y\|^{2}$, and $D_{\textup{Euc}}=\sqrt{2}D$, we obtain
$f(x_{t})-f^{*}\leq\frac{2\mu LD_{\textup{Euc}}^{2}}{\rho(t+2)},$
Theorem 5.5 requires $\rho/\sigma\leq L$ and $\frac{3\rho}{\mu}\leq 2-\nu$. These assumptions are easy to satisfy.
Example 5.6 (Example 3.1 in the arXiv version of [Liao2024-vd]).
Let us consider
$f(x)=\begin{cases}-x^{2}+1,&\text{if}\ -1<x<-0.5,\\ 3(x+1)^{2},&\text{otherwise}.\end{cases}$
The function $f$ is not convex but $\rho$-weakly convex with $\rho=2$. A global optimal solution of $f$ is $x^{*}=-1$ and its value is $f(x^{*})=0$. Moreover, $f$ has the quadratic growth property with $0<\mu\leq 6$. Because $f$ is a quadratic function, we have $L\geq 6$. It holds that $2=\rho\leq L$ and $1=\frac{3\rho}{\mu}\leq 1<2-\nu$ (set $\mu=6$). Therefore, the assumption of Theorem 5.5 holds.
In order to verify $\frac{3\rho}{\mu}\leq 2-\nu$, it is easier to examine the sufficient condition $\frac{3\rho}{\mu}\leq 1$ instead because the exact value of $\nu$ is difficult to estimate. On the other hand, without loss of generality, we can assume $\sigma=1$. When $\sigma>1$, $\rho/\sigma<\rho\leq L$ holds. When $\sigma<1$, we can use $\phi_{1}=\phi+\frac{1-\sigma}{2}\|\cdot\|^{2}$, which is 1-strongly convex. When $\phi$ is convex but not strongly convex, we can use $\phi_{2}=\phi+\frac{1}{2}\|\cdot\|^{2}$. Therefore, it suffices to verify $\rho/\sigma\leq\rho\leq L$, which often holds (for example, see an example of phase retrieval in Section 6.3). Next we establish local accelerated convergence with the short step step-size, i.e., $\gamma_{t}=\min\left\{\left(\frac{\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle}{L(1+\nu)D_{\phi}(v_{t},x_{t})}\right)^{\frac{1}{\nu}},1\right\}$.
Theorem 5.7 (Local accelerated convergence for inner optima).
Suppose that Assumptions 3.1 and 5.2 hold. Let $D:=\sqrt{\sup_{x,y\in P}D_{\phi}(x,y)}$ be the diameter. Assume that there exists a minimizer $x^{*}\in\operatorname{int}P$, i.e., there exists an $r>0$ with $B(x^{*},r)\subset P$. Consider the iterates of Algorithm 1 with $\gamma_{t}=\min\left\{\left(\frac{\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle}{L(1+\nu)D_{\phi}(v_{t},x_{t})}\right)^{\frac{1}{\nu}},1\right\}$. Then, if $\rho/\mu<1$, it holds that, for all $t\geq 1$,
$f(x_{t})-f^{*}\leq\begin{cases}\max\left\{\frac{1}{2}\left(1+\frac{\rho}{\mu}\right),1-\frac{r^{2}}{2Mc^{2}D^{2}}\right\}^{t-1}LD^{2}&\text{if}\ \nu=1,\\ \left(\frac{1}{1+\nu}\left(1+\frac{\nu\rho}{\mu}\right)\right)^{t-1}LD^{2}&\text{if}\ 1\leq t\leq t_{0},\nu\in(0,1),\\ \frac{(L(1+\nu)(1-\rho/\mu)^{\nu}c^{1+\nu}D^{2}/r^{1+\nu})^{2/(1-\nu)}}{\left(1+\frac{1-\nu}{2(1+\nu)}\left(1-\frac{\rho}{\mu}\right)(t-t_{0})\right)^{2\nu/(1-\nu)}}=\mathcal{O}(1/t^{2\nu/(1-\nu)})&\text{if}\ t\geq t_{0},\nu\in(0,1),\end{cases}$
where $c=\frac{\sqrt{2\mu}}{\mu-\rho}$, $M:=(L(1+\nu))^{1/\nu}$, and
$t_{0}:=\max\left\{\left\lfloor\log_{\frac{1}{1+\nu}\left(1+\frac{\nu\rho}{\mu}\right)}\left(\frac{(L(1+\nu)(1-\rho/\mu)^{\nu}c^{1+\nu}D^{2}/r^{1+\nu})^{2/(1-\nu)}}{LD^{2}}\right)\right\rfloor+2,1\right\}.$
Proof.
Let $h_{t}:=f(x_{t})-f^{*}$. Using Lemma 3.3, we have
$h_{t}-h_{t+1}=f(x_{t})-f(x_{t+1})\geq\frac{\nu}{1+\nu}\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle\gamma_{t},$
where $\gamma_{t}=\min\left\{1,\left(\frac{\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle}{L(1+\nu)D_{\phi}(v_{t},x_{t})}\right)^{\frac{1}{\nu}}\right\}$. We consider two cases: (i) $\gamma_{t}<1$ and (ii) $\gamma_{t}=1$.
(i) $\gamma_{t}<1$: We have
$h_{t}-h_{t+1}$ $\geq\frac{\nu\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle}{1+\nu}\min\left\{1,\left(\frac{\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle}{L(1+\nu)D_{\phi}(v_{t},x_{t})}\right)^{1/\nu}\right\}$
$\geq\frac{\nu}{1+\nu}\frac{\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle^{1+1/\nu}}{(L(1+\nu)D_{\phi}(v_{t},x_{t}))^{1/\nu}}$
$\geq\frac{\nu}{1+\nu}\frac{r^{1+1/\nu}\|\nabla f(x_{t})\|^{1+1/\nu}}{(L(1+\nu))^{1/\nu}D^{2/\nu}}$
$\geq\frac{\nu}{1+\nu}\frac{r^{1+1/\nu}}{Mc^{1+1/\nu}D^{2/\nu}}h_{t}^{\frac{1+\nu}{2\nu}},$
where the third inequality holds due to Proposition 2.17 and the definition of $D$ and the last inequality holds due to the local PL inequality from Remark 2.15 with $c=\frac{\sqrt{2\mu}}{\mu-\rho}$ and $M:=(L(1+\nu))^{1/\nu}$.
(ii) In the case where $\gamma_{t}=1$, i.e., $\left(\frac{\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle}{L(1+\nu)D_{\phi}(v_{t},x_{t})}\right)^{\frac{1}{\nu}}\geq 1$, this implies
$\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle\geq L(1+\nu)D_{\phi}(v_{t},x_{t}),$ (5.7)
because of $11/\nu>$ 1. Using Lemma 2.8 and (5.7), we have
$h_{t+1}-h_{t}$ $\leq LD_{\phi}(v_{t},x_{t})-\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle$
$\leq-\frac{\nu}{1+\nu}\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle$
$\leq-\frac{\nu}{1+\nu}\left(h_{t}-\frac{\rho}{2}\|x_{t}-x^{*}\|^{2}\right)$
$\leq-\frac{\nu}{1+\nu}\left(1-\frac{\rho}{\mu}\right)h_{t},$
where the third inequality holds because of (5.1) in Lemma 5.3, and the last inequality holds because of the local quadratic growth condition of $f$ with $\mu>0$. Therefore, we have
$h_{t+1}\leq\frac{1}{1+\nu}\left(1+\frac{\rho\nu}{\mu}\right)h_{t}.$
From (i) and (ii), we have
$h_{t}-h_{t+1}\geq\min\left\{\frac{\nu}{1+\nu}\left(1-\frac{\rho}{\mu}\right)h_{t},\frac{\nu}{1+\nu}\frac{r^{1+1/\nu}}{Mc^{1+1/\nu}D^{2/\nu}}h_{t}^{\frac{1+\nu}{2\nu}}\right\}.$
When $\nu=1$, we have $h_{t}-h_{t+1}\geq\min\left\{\frac{1}{2}\left(1-\frac{\rho}{\mu}\right),\frac{r^{2}}{2Mc^{2}D^{2}}\right\}\cdot h_{t}$. This inequality and the initial bound $f(x_{1})-f^{*}\leq LD^{2}$ due to Lemma 2.7 imply
$h_{t}\leq\max\left\{\frac{1}{2}\left(1+\frac{\rho}{\mu}\right),1-\frac{r^{2}}{2Mc^{2}D^{2}}\right\}^{t-1}LD^{2}.$
On the other hand, when $\nu\in(0,1)$,
$h_{t}-h_{t+1}\geq\min\left\{\frac{\nu}{1+\nu}\left(1-\frac{\rho}{\mu}\right),\frac{\nu}{1+\nu}\frac{r^{1+1/\nu}}{Mc^{1+1/\nu}D^{2/\nu}}h_{t}^{\frac{1-\nu}{2\nu}}\right\}\cdot h_{t}.$
Using $f(x_{1})-f^{*}\leq LD^{2}$ and Lemma 2.20 with $c_{0}=LD^{2}$, $c_{1}=\frac{\nu}{1+\nu}\left(1-\frac{\rho}{\mu}\right)$, $c_{2}=\frac{\nu}{1+\nu}\frac{r^{1+1/\nu}}{Mc^{1+1/\nu}D^{2/\nu}}$, and $\theta_{0}=\frac{1-\nu}{2\nu}>0$, we have the claim. ∎
When $\phi=\frac{1}{2}\|\cdot\|^{2}$ and $\nu=1$, we have local linear convergence with $\gamma_{t}=\min\left\{\frac{\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle}{L\|v_{t}-x_{t}\|^{2}},1\right\}$ from Theorem 5.7. When $\phi=\frac{1}{2}\|\cdot\|$, i.e., $D_{\phi}(x,y)=\frac{1}{2}\|x-y\|^{2}$, $D_{\textup{Euc}}=\sqrt{2}D$ provides
$f(x_{t})-f^{*}\leq\max\left\{\frac{1}{2}\left(1+\frac{\rho}{\mu}\right),1-\frac{r^{2}}{2Lc^{2}D_{\textup{Euc}}^{2}}\right\}^{t-1}\frac{LD_{\textup{Euc}}^{2}}{2}.$
Because Algorithm 2 is well-defined (see also Remark 3.4), the convergence result of the FW algorithm with Algorithm 2 is the same as Theorem 5.7.
Finally, we establish the local linear convergence of Algorithm 3. In the same way as Theorem 4.6, the pyramidal width is the minimal $\delta>0$ satisfying Lemma 2.18 (see also [Braun2022-ew, Lemma 2.26] or [Lacoste-Julien2015-gb, Theorem 3]). An upper bound exists on the primal gap for weakly convex functions. Its proof can be found in Appendix A.1.
Lemma 5.8 (Upper bound on primal gap for weakly convex functions).
Suppose that Assumptions 3.1 and 5.2 hold. Let $P$ be a polytope with the pyramidal width $\delta>0$. Let $\mathcal{S}$ denote any set of vertices of $P$ with $x\in\operatorname{conv}\mathcal{S}$. Let $\psi$ be any vector, so that we define $v^{\textup{FW}}=\operatorname*{argmin}_{v\in P}\langle\psi,v\rangle$ and $v^{\textup{A}}=\operatorname*{argmax}_{v\in\mathcal{S}}\langle\psi,v\rangle$. If $\rho/\mu<1$, then it holds that, for all $x\in[f\leq f^{*}+\zeta]\cap P$,
$f(x)-f^{*}\leq\frac{2\mu}{(\mu-\rho)^{2}\delta^{2}}\langle\nabla f(x),v^{\textup{A}}-v^{\textup{FW}}\rangle^{2}.$ (5.8)
We are ready to prove the local linear convergence of Algorithm 3 for nonconvex optimization. We also assume $\rho\leq L$, which is not restrictive (see also the discussion of Theorem 5.5 and Example 5.6).
Theorem 5.9 (Local linear convergence by the away-step FW algorithm).
Suppose that Assumptions 3.14.5, and 5.2 hold. Let $P\subset\mathbb{R}^{n}$ be a polytope. The convergence rate of Algorithm 3 with $f$ is linear: if $\rho<\mu\leq L$, for all $t\geq 1$
$f(x_{t})-f^{*}\leq\begin{cases}2\left(1-\frac{\omega}{4L}\frac{\delta^{2}}{D^{2}}\right)^{\lceil(t-1)/2\rceil}LD^{2}&\text{if}\ \nu=1,\\ 2\left(\frac{1}{1+\nu}\left(1+\frac{\nu\rho}{\mu}\right)\right)^{\lceil(t-1)/2\rceil}LD^{2}&\text{if}\ 1\leq t\leq t_{0},\nu\in(0,1),\\ \frac{(L^{2}(1+\nu)^{2}D^{4}(1-\rho/\mu)^{2\nu}/(\omega\delta^{2})^{1+\nu})^{1/(1-\nu)}}{\left(1+\frac{1-\nu}{2(1+\nu)}\left(1-\frac{\rho}{\mu}\right)\lceil(t-t_{0})/2\rceil\right)^{2\nu/(1-\nu)}}=\mathcal{O}(1/t^{2\nu/(1-\nu)})&\text{if}\ t\geq t_{0},\nu\in(0,1),\end{cases}$
where $D:=\sqrt{\sup_{x,y\in P}D_{\phi}(x,y)}$ and $\delta$ are the diameter and the pyramidal width of the polytope $P$, respectively, and $\omega:=\frac{(\mu-\rho)^{2}}{8\mu}$, and
$t_{0}:=\max\left\{\left\lfloor\log_{\frac{1}{1+\nu}\left(1+\frac{\nu\rho}{\mu}\right)}\frac{(L^{2}(1+\nu)^{2}D^{4}(1-\rho/\mu)^{2\nu}/(\omega\delta^{2})^{1+\nu})^{1/(1-\nu)}}{2LD^{2}}\right\rfloor+2,1\right\}.$
Proof.
By letting $h_{t}=f(x_{t})-f^{*}$ and using Lemma 5.8, we have
$h_{t}\leq\frac{2\mu}{(\mu-\rho)^{2}\delta^{2}}\langle\nabla f(x_{t}),v^{\textup{A}}-v^{\textup{FW}}\rangle^{2}\leq\frac{8\mu\langle\nabla f(x_{t}),d_{t}\rangle^{2}}{(\mu-\rho)^{2}\delta^{2}}=\frac{\langle\nabla f(x_{t}),d_{t}\rangle^{2}}{\omega\delta^{2}},$ (5.9)
where the second inequality holds because $d_{t}$ is either $x_{t}-v_{t}^{\textup{FW}}$ or $v_{t}^{\textup{A}}-x_{t}$ with $\langle\nabla f(x_{t}),d_{t}\rangle\geq\langle\nabla f(x),v^{\textup{A}}-v^{\textup{FW}}\rangle/2$ in Lines 3 and 3 of Algorithm 3. From Lemma 3.3 with $\gamma_{t,\max}=1$ (Frank–Wolfe steps) and $\gamma_{t,\max}=\frac{\lambda_{v_{t}^{\textup{A}},t}}{1-\lambda_{v_{t}^{\textup{A}},t}}$ (away steps), we obtain
$h_{t}-h_{t+1}$ $\geq\frac{\nu}{1+\nu}\langle\nabla f(x_{t}),d_{t}\rangle\min\left\{\gamma_{t,\max},\left(\frac{\langle\nabla f(x_{t}),d_{t}\rangle}{L(1+\nu)D_{\phi}(v_{t},x_{t})}\right)^{\frac{1}{\nu}}\right\}$
$=\min\left\{\gamma_{t,\max}\frac{\nu}{1+\nu}\langle\nabla f(x_{t}),d_{t}\rangle,\frac{\nu}{1+\nu}\frac{\langle\nabla f(x_{t}),d_{t}\rangle^{1+1/\nu}}{(L(1+\nu)D_{\phi}(v_{t},x_{t}))^{1/\nu}}\right\}$
$\geq\min\left\{\frac{\nu\gamma_{t,\max}}{1+\nu}\left(1-\frac{\rho}{\mu}\right)h_{t},\frac{\nu}{1+\nu}\frac{(h_{t}\omega\delta^{2})^{\frac{1+\nu}{2\nu}}}{(L(1+\nu)D^{2})^{1/\nu}}\right\}$
$=\min\left\{\frac{\nu\gamma_{t,\max}}{1+\nu}\left(1-\frac{\rho}{\mu}\right)h_{t},\frac{\nu}{1+\nu}\frac{(\omega\delta^{2})^{\frac{1+\nu}{2\nu}}}{(L(1+\nu)D^{2})^{1/\nu}}h_{t}^{\frac{1+\nu}{2\nu}}\right\},$
where the second inequality holds because of $\langle\nabla f(x_{t}),d_{t}\rangle\geq\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle\geq h_{t}-\frac{\rho}{2}\|x_{t}-v_{t}\|^{2}\geq\left(1-\frac{\rho}{\mu}\right)h_{t}$ (from Lemma 5.3 and the quadratic growth condition of $f$) and (5.9).
In $\nu=1$, for Frank–Wolfe steps, we have $\gamma_{t,\max}\left(1-\frac{\rho}{\mu}\right)=\left(1-\frac{\rho}{\mu}\right)\geq\frac{1}{8}\left(1-\frac{\rho}{\mu}\right)^{2}\frac{\delta^{2}}{D^{2}}=\frac{\omega\delta^{2}}{\mu D^{2}}\geq\frac{\omega\delta^{2}}{2LD^{2}}$ by $0<\rho<\mu\leq L$ and $\delta\leq D$. For away steps, it seems that we only obtain a monotone progress $h_{t+1}<h_{t}$ because it is difficult to estimate $\gamma_{t,\max}$ below. However, $\gamma_{t,\max}$ cannot be small too often. Let us consider $\gamma_{t}=\gamma_{t,\max}$ in an away step. In that case, $v^{\textup{A}}_{t}$ is removed from the active set $\mathcal{S}_{t+1}$ in line 3. Moreover, the active set can only grow in Frank–Wolfe steps (line 3). It is impossible to remove more vertices from $\mathcal{S}_{t+1}$ than have been added in Frank–Wolfe steps. Therefore, at most half of iterations until $t$ iterations are in away steps, i.e., in all other steps, we have $\gamma_{t}=\frac{\langle\nabla f(x_{t}),d_{t}\rangle}{2LD_{\phi}(v_{t},x_{t})}<\gamma_{t,\max}$ and then $h_{t+1}\leq\left(1-\frac{\omega\delta^{2}}{4LD^{2}}\right)h_{t}$. Because we have $h_{t+1}\leq\left(1-\frac{\omega\delta^{2}}{4LD^{2}}\right)h_{t}$ for at least half of the iterations and $h_{t+1}\leq h_{t}$ for the rest, using $h_{1}\leq 2LD^{2}$ from (5.6) and $\rho<L$, we obtain
$f(x_{t})-f^{*}\leq 2\left(1-\frac{\omega}{4L}\frac{\delta^{2}}{D^{2}}\right)^{\lceil(t-1)/2\rceil}LD^{2}.$
In $\nu\in(0,1)$, we have $h_{t}-h_{t+1}\geq\min\left\{\frac{\nu}{1+\nu}\left(1-\frac{\rho}{\mu}\right),\frac{\nu}{1+\nu}\frac{(\omega\delta^{2})^{\frac{1+\nu}{2\nu}}}{(L(1+\nu)D^{2})^{1/\nu}}h_{t}^{\frac{1-\nu}{2\nu}}\right\}\cdot h_{t}$ for at least half of the iterations and $h_{t+1}\leq h_{t}$ for the rest. The initial bound $f(x_{1})-f^{*}\leq(\rho+L)D^{2}\leq 2LD^{2}$ holds from (5.6) and $\rho<L$. We use Lemma 2.20 with $c_{0}=2LD^{2}$, $c_{1}=\frac{\nu}{1+\nu}\left(1-\frac{\rho}{\mu}\right)$, $c_{2}=\frac{\nu}{1+\nu}\frac{(\omega\delta^{2})^{\frac{1+\nu}{2\nu}}}{(L(1+\nu)D^{2})^{1/\nu}}$, and $\theta_{0}=\frac{1-\nu}{2\nu}$ and obtain the claim. ∎
When $\phi=\frac{1}{2}\|\cdot\|^{2}$ and $\nu=1$, the local linear convergence of Algorithm 3 is equivalent to
$f(x_{t})-f(x^{*})\leq\left(1-\frac{\omega}{2L}\frac{\delta^{2}}{D_{\textup{Euc}}^{2}}\right)^{\lceil(t-1)/2\rceil}LD_{\textup{Euc}}^{2}.$
Without loss of generality, we set $\sigma=1$. If $\rho\leq L$ does not hold, we can use $h_{1}\leq(\rho+L)D^{2}$ as the initial bound.
Moreover, when $P$ is $\alpha$-strongly convex set and $\phi=\frac{1}{2}\|\cdot\|^{2}$, we also establish local linear convergence. Its proof can be found in Appendix A.2.
6 Numerical Experiments
In this section, we conducted numerical experiments to examine the performance of our proposed algorithms. All numerical experiments were performed in Julia 1.11 using the FrankWolfe.jl package [Besancon2022-lx] [^2] https://github.com/ZIB-IOL/FrankWolfe.jl on a MacBook Pro with an Apple M2 Max and 64GB LPDDR5 memory.
We compare our algorithm with the following algorithms and use the following notation:
  • BregFW: the FW algorithm with the adaptive Bregman step-size strategy (Algorithm 2, our proposed)
  • BregAFW: the away-step FW algorithm with the adaptive Bregman step-size strategy (Algorithm 3 using Algorithm 2, i.e., $L_{t},\nu_{t},\gamma_{t}\leftarrow\textnormal{{step\_size}}(f,\phi,x_{t},v_{t},L_{t-1},\gamma_{t,\max})$, our proposed update)
  • EucFW: the FW algorithm with the adaptive (Euclidean) step-size strategy [Pedregosa2020-ay]
  • EucAFW: the away-step FW algorithm with the adaptive (Euclidean) step-size strategy [Pedregosa2020-ay]
  • ShortFW: the FW algorithm with the (Euclidean) short step
  • ShortAFW: the away-step FW algorithm with the (Euclidean) short step
  • OpenFW: the FW algorithm with the open loop with $\gamma_{t}=\frac{2}{2+t}$
  • OpenAFW: the away-step FW algorithm with the open loop with $\gamma_{t}=\frac{2}{2+t}$
  • MD: the mirror descent [Nemirovskij1983-sm]
Note that we included OpenAFW only for comparison purposes; there is no established convergence theory for AFW with open-loop strategies. In particular, there are no proper drop steps, and the favorable properties of the away-step Frank-Wolfe algorithm are lost; see [Braun2022-ew]. We use $\beta=0.9$, $\eta=0.9$, $\tau=2$, and $\gamma_{\max}=1$ throughout all numerical experiments.
6.1 Nonnegative Linear Inverse Problems
Given a nonnegative matrix $A\in\mathbb{R}_{+}^{m\times n}$ and a nonnegative vector $b\in\mathbb{R}_{+}^{m}$, the goal of nonnegative (Poisson) linear inverse problems is to recover a signal $x\in\mathbb{R}_{+}^{n}$ such that $Ax\simeq b$. This class of problems has been studied in image deblurring [Bertero2009-jq] and positron emission tomography [Vardi1985-de] as well as optimization [Bauschke2017-hg, Takahashi2024-wz]. Since the dimension of $x$ is often larger than the number of observations $m$, the system is indeterminate. From this point of view, we consider the constraint $\Delta_{n}:=\{x\in\mathbb{R}_{+}^{n}\,|\,\sum_{j=1}^{n}x_{j}\leq 1\}$. Recovering $x$ can be formulated as a minimization problem:
$\min_{x\in\Delta_{n}}\quad f(x):=d(Ax,b),$ (6.1)
where $d(x,y):=\sum_{i=1}^{m}\left(x_{i}\log\frac{x_{i}}{y_{i}}+y_{i}-x_{i}\right)$ is the Kullback–Leibler divergence (see also Example 2.3). Problem (6.1) is convex, while $\nabla f$ is not Lipschitz continuous on $\mathbb{R}_{+}^{n}$. The pair $(f,\phi)$ is $L$-smad on $\mathbb{R}_{+}^{n}$ with $\phi(x)=\sum_{j=1}^{n}x_{j}\log x_{j}$ and $L\geq\max_{1\leq j\leq n}\sum_{i=1}^{m}a_{ij}$ from [Bauschke2017-hg, Lemma 8].
We compared BregFW with EucFW, ShortFW, OpenFW, and the mirror descent algorithm (MD[Nemirovskij1983-sm]. The subproblem of MD can be solved in closed-form for $\{x\in\mathbb{R}_{+}^{n}\,|\,\sum_{j=1}^{n}x_{j}=1\}$ by [Beck2017-qc, Example 3.71], and this can be readily extended to $\Delta_{n}$. We used $1000$ as maximum iteration limit and we generated $\tilde{A}$ from an i.i.d. normal distribution and set $a_{ij}=|\tilde{a}_{ij}|/\sum_{i=1}^{m}|\tilde{a}_{ij}|$. We also generated $\tilde{x}$ from an i.i.d. uniform distribution in $[0,1]$ and set the ground truth $x^{*}=0.8\tilde{x}/\sum_{j=1}^{n}\tilde{x}_{j}$ so that $x^{*}\in\operatorname{int}\Delta_{n}$. All components of the initial point $x_{0}$ were $1/n$. For $(m,n)=(100,1000)$, Figure 1 shows the primal gap $f(x_{t})-f^{*}$ and the FW gap $\max_{v\in P}\langle\nabla f(x_{t}),x_{t}-v\rangle$ per iteration (left) and the primal gap per second (right). Table 2 shows average numbers of the primal gap, the FW gap, and computational time over 20 different instances for $(m,n)=(100,1000)$. Here, BregFW outperformed other algorithms, both in iterations and time.
Figure 1: Nonnegative linear inverse problem for $(m,n)=(100,1000)$.
Table 2: Average numbers of the primal gap, the FW gap, and computational time (s) over 20 different instances of nonnegative linear inverse problems for $(m,n)=(100,1000)$.
algorithm primal gap FW gap time (s)
BregFW 6.963691e-08 1.145520e-05 1.457719e-01
EucFW 3.028696e-07 2.922331e-05 1.077826e-01
ShortFW 4.747044e-05 4.613479e-04 6.251837e-02
OpenFW 4.957628e-07 8.538245e-05 6.503594e-02
MD 2.249368e-06 1.721359e-04 1.888499e-01
6.2 $\ell_{p}$oss Problem
We consider the $\ell_{p}$ loss problem [Kyng2015-uo, Maddison2021-nt] to find $x\in P$ such that $Ax\simeq b$, defined by
$\min_{x\in P}\quad\|Ax-b\|_{p}^{p},$ (6.2)
where $A\in\mathbb{R}^{m\times n}$, $b\in\mathbb{R}^{m}$, and $p>1$. Defining $f(x)=\|Ax-b\|_{p}^{p}$, we have
$\nabla f(x)$ $=p\sum_{i=1}^{m}|\langle a_{i},x\rangle-b_{i}|^{p-1}\operatorname{sgn}(\langle a_{i},x\rangle-b_{i}),$
$\nabla^{2}f(x)$ $=p(p-1)\sum_{i=1}^{m}|\langle a_{i},x\rangle-b_{i}|^{p-2}a_{i}a_{i}^{\mathsf{T}},$
where $a_{i}$ is the $i$th row vector of $A$. Problem (6.2) is convex, but $\nabla f$ is not Lipschitz continuous on $\mathbb{R}^{n}$ when $p\neq 2$. Furthermore, when $p<2$, $f$ is not $\mathcal{C}^{2}$ but $\mathcal{C}^{1}$. Therefore, for $21<p<$ 2, $\nabla f$ is also not Lipschitz continuous over compact sets. Since $f$ is convex, the pair $(f,\phi)$ is 1-smad with $\phi=f$.
We use an $\ell_{2}$ norm ball as a constraint, i.e., $P=\{x\in\mathbb{R}^{n}\,|\,\|x\|_{2}\leq 1\}$. We compared BregFW with EucFW, ShortFW, and OpenFW. We generated $A$ from an i.i.d. normal distribution and normalized it so that $\|a_{i}\|=1$. We also generated $\tilde{x}$ from an i.i.d. normal distribution and set $x^{*}=0.8\tilde{x}/\|\tilde{x}\|$ to ensure $x^{*}\in\operatorname{int}P$. The initial point $x_{0}$ was generated by computing an extreme point of $P$ that minimizes the linear approximation of $f$. For $(n,m)=(100,100)$ and $p=1.1$, Figure 2 shows the primal gap $f(x_{t})-f^{*}$ and the FW gap $\max_{v\in P}\langle\nabla f(x_{t}),x_{t}-v\rangle$ per iteration and those gaps per second up to 1000 iterations. Table 3 also shows the average performance over 20 different instances. BregFW outperformed the other algorithms. Since $\nabla f$ is not Lipschitz continuous, ShortFW did not reduce the primal and FW gap.
Figure 2: The $\ell_{p}$ loss problem for $(m,n)=(1000,100)$.
Table 3: Average numbers of the primal gap, FW gap, and computational time (s) over 20 different instances of $\ell_{p}$ loss problems for $(m,n)=(1000,100)$.
algorithm primal gap FW gap time (s)
BregFW 1.056988e-13 3.764084e-01 5.600876e-01
EucFW 6.341301e-10 8.576574e-01 3.138045e-01
ShortFW 2.465054e+01 4.754413e+01 9.198800e-02
OpenFW 1.698968e-02 4.007941e+00 8.713342e-02
6.3 Phase Retrieval
We are interested in phase retrieval, which involves finding a signal $x\in\mathbb{R}^{n}$ such that
$|\langle a_{i},x\rangle|^{2}\simeq b_{i},\quad i=1,\ldots,m,$
where $a_{i}\in\mathbb{R}^{n}$ describes the model and $b\in\mathbb{R}^{m}$ is a vector of measurements. Phase retrieval has a long history. Patterson studied phase retrieval in X-ray crystallography in 1934 [Patterson1934-sg] and 1944 [Patterson1944-nm]. Phase retrieval arises in many fields of science and engineering, such as image processing [Candes2015-mk], X-ray crystallography [Patterson1934-sg, Patterson1944-nm], and optics [Shechtman2015-ic]. Bolte et al. [Bolte2018-zt] applied the Bregman proximal gradient algorithm to phase retrieval.
In applications, $a_{i}$ and $x$ are often complex vectors. Now, we consider real vectors. To achieve the goal of phase retrieval, we focus on the following nonconvex optimization problem:
$\min_{x\in P}\quad f(x):=\frac{1}{4}\sum_{i=1}^{m}\left(|\langle a_{i},x\rangle|^{2}-b_{i}\right)^{2},$
where $P\subset\mathbb{R}^{n}$ is a compact convex set. We introduce some properties of $f$. Let $\phi$ be defined by
$\phi(x)=\frac{1}{4}\|x\|^{4}+\frac{1}{2}\|x\|^{2},$
which is $\sigma$-strongly convex for $\sigma\leq 1$. For any $L$ satisfying
$L\geq\sum_{i=1}^{m}\left(3\|a_{i}\|^{4}+\|a_{i}\|^{2}|b_{i}|\right),$ (6.3)
the pair $(f,\phi)$ is $L$-smad on $\mathbb{R}^{n}$ [Bolte2018-zt, Lemma 5.1]. The right-hand side of (6.3) is relatively large in applications, while Takahashi et al. derived a smaller lower bound of $L$ [Takahashi2023-uh, Propositions 5 and 6]. In addition, because we have
$\nabla^{2}f(x)=\sum_{i=1}^{m}\left((3|\langle a_{i},x\rangle|^{2}-b_{i})a_{i}a_{i}^{\mathsf{T}}\right),$
the matrix $\nabla^{2}f(x)+\rho I$ is positive semidefinite for $\rho\geq\sum_{i=1}^{m}\|a_{i}\|^{2}|b_{i}|$, i.e., $f$ is $\rho$-weakly convex. For example, $L=\sum_{i=1}^{m}\left(3\|a_{i}\|^{4}+\|a_{i}\|^{2}|b_{i}|\right)\geq\sum_{i=1}^{m}\|a_{i}\|^{2}|b_{i}|=\rho/\sigma$ holds with $\sigma=1$.
We use a $K$-sparse polytope as a constraint, i.e., $P=\{x\in\mathbb{R}^{n}\,|\,\|x\|_{1}\leq K,\|x\|_{\infty}\leq 1\}$. We compared BregFW with EucFW, ShortFW, and OpenFW. The maximum iteration is 1000. We generated $a_{i}$, $i=1,\ldots,m$ from an i.i.d. normal distribution and normalized it to have norm 1. We also generated $x^{*}$ from an i.i.d. uniform distribution in $[0,1]$ and normalized $x^{*}$ to have sum 1. The initial point $x_{0}$ was generated by computing an extreme point of $P$ that minimizes the linear approximation of $f$. For $(m,n)=(100,2000)$ and $K=200$, Figure 3 shows the primal and FW gaps per iteration and gaps per second. Because $\nabla f$ is not Lipschitz continuous, ShortFW does not converge, and EucFW performs slowly. The primal gap and the FW gap by the Bregman adaptive step-size strategy are the smallest among these step-size strategies. Table 4 shows the average performance over 20 different instances of $(m,n)=(100,2000)$. We compared BregAFW with EucAFW, ShortAFW, and OpenAFW. In only this setting, we generated $x^{*}$ from an i.i.d. uniform distribution in $[0,1]$ and did not normalize it; that is, $x^{*}$ might be in the face of $P$. Figure 4 shows another setting’s results for $(m,n)=(200,200)$ and $K=110$. Table 5 shows the average performance over 20 different instances for $(m,n)=(200,200)$ and $K=110$. The primal gap by BregAFW is the smallest among these algorithms, while ShortAFW has the smallest value of the FW gap.
Figure 3: Phase retrieval for $(m,n)=(100,2000)$ and $K=200$.
Table 4: Average numbers of the primal gap, FW gap, and computational time (s) over 20 different instances of phase retrieval for $(m,n)=(100,2000)$ and $K=200$.
algorithm primal gap FW gap time
BregFW 3.307714e-09 4.372800e-06 9.532406e-01
EucFW 1.419221e-04 1.271594e+01 8.291155e-01
ShortFW 5.488362e+04 5.389255e+03 1.822046e-01
OpenFW 4.710460e-08 2.747374e+00 1.945798e-01
Figure 4: Phase retrieval for $(m,n)=(200,200)$ and $K=110$ via away-step FW algorithms.
Table 5: Average numbers of the primal gap, FW gap, and computational time (sec) over 20 different instances of phase retrieval for $(m,n)=(200,200)$ and $K=110$ via away-step FW algorithms.
algorithm primal gap FW gap time (s)
BregAFW 4.343027e+00 9.757773e-01 2.333602e-01
EucAFW 4.396803e+00 1.221226e+00 3.942057e-01
ShortAFW 4.409245e+00 2.597756e-02 7.353968e-02
OpenAFW 4.409352e+00 7.199201e-02 5.899826e-02
6.4 Low-Rank Minimization
Given a symmetric matrix $M\in\mathbb{R}^{n\times n}$, our goal is to find $X\in\mathbb{R}^{n\times r}$ such that $M\simeq XX^{\mathsf{T}}$. This is accomplished by minimizing the function
$\min_{X\in P}\quad f(X):=\frac{1}{2}\|XX^{\mathsf{T}}-M\|_{F}^{2},$
where $P\subset\mathbb{R}^{n\times r}$. We assume that $r\leq n$. This problem is known as low-rank minimization [Dragomir2021-rv]. In this paper, we define $P=\{X\in\mathbb{R}^{n\times r}\,|\,\|X\|_{*}\leq\xi\}$, where $\|\cdot\|_{*}$ denotes the nuclear norm and $r\in\mathbb{R}_{+}$ for the low-rank assumption.
We define
$\phi(X)=\frac{1}{4}\|X\|_{F}^{4}+\frac{1}{2}\|X\|_{F}^{2}.$
There exists a constant $L$ such that the pair $(f,\phi)$ is $L$-smad on $\mathbb{R}^{n}$ [Dragomir2021-rv]. Additionally, $f(X)$ is weakly convex on any compact set due to Proposition 2.10, which follows from the twice continuous differentiability of $f$.
We also compared BregFW with EucFW, ShortFW, and OpenFW. The parameter settings are the same as those in the previous subsection. We generated $X^{*}$ from an i.i.d. uniform distribution in $[0,1]$, normalized each column of $X^{*}$, and set $M=X^{*}(X^{*})^{\mathsf{T}}$. The initial point $X_{0}$ was generated from an i.i.d. uniform distribution in $[0,1]$. We set $\xi=10\lambda_{\max}(M)$ for $P$. Figure 5 shows the primal and FW gaps per iteration and gaps per second for $(n,r)=(1000,20)$ up to the 1000th iteration. Table 6 presents the average performance over 20 different instances for $(n,r)=(1000,20)$. BregFW performs slightly better than EucFW. OpenFW also performed as fast as BregFW and EucFW, but its performance was unstable. ShortFW did not converge due to the lack of Lipschitz continuity of $\nabla f$.
Figure 5: Low-rank minimization for $(n,r)=(1000,20)$.
Table 6: Average values of the primal gap, FW gap, and computational time (s) over 20 different instances of low-rank minimization for $(n,r)=(1000,20)$.
algorithm primal gap FW gap time (s)
BregFW 5.651947e-01 1.472666e+01 1.260218e+01
EucFW 5.776476e-01 1.723729e+01 1.228013e+01
ShortFW 2.710108e+08 2.169351e+09 4.137911e+00
OpenFW 1.978179e+00 2.121183e+03 4.176979e+00
6.5 Nonnegative Matrix Factorization
Given a nonnegative matrix $V\in\mathbb{R}_{+}^{m\times n}$, nonnegative matrix factorization (NMF) aims to find nonnegative matrices $W\in\mathbb{R}_{+}^{m\times r}$ and $H\in\mathbb{R}_{+}^{r\times n}$ such that $V\simeq WH$. NMF can be formulated as a minimization problem of the loss function that measures the difference between $V$ and $WH$, i.e.,
$\min_{(W,H)\in P}\quad f(W,H):=\frac{1}{2}\|WH-V\|_{F}^{2},$ (6.4)
where $P$ is a compact convex subset of $\mathbb{R}_{+}^{m\times r}\times\mathbb{R}_{+}^{r\times n}$. The objective function $f$ is weakly convex over $P$ due to Proposition 2.10. The gradient $\nabla f$ is not Lipschitz continuous, while $(f,\phi)$ is smooth adaptable [Mukkamala2019-mk] with $\phi(W,H)=\frac{1}{4}(\|W\|_{F}^{2}+\|H\|_{F}^{2})^{2}+\frac{1}{2}(\|W\|_{F}^{2}+\|H\|_{F}^{2})$.
[…]
Figure 6: NMF for $(m,n,r)=(100,5000,20)$.
Table 7: Average values of the primal gap, FW gap, and computational time (s) over 20 different instances of NMF for $(m,n,r)=(100,5000,20)$.
algorithm primal gap FW gap time (s)
BregFW 1.201735e-04 1.151457e-02 5.098349e+01
EucFW 1.281873e-04 7.625941e-04 4.359247e+01
[…]
Figure 7: NMF for MovieLens 100K Dataset.
Acknowledgement
This project has been partially supported by the Japan Society for the Promotion of Science (JSPS) through Invitational Fellowships for Research in Japan program (L24503), JSPS KAKENHI Grant Number JP23K19953 and JP25K21156; JSPS KAKENHI Grant Number JP23H03351 and JST ERATO Grant Number JPMJER1903. Research reported in this paper was also partially supported the Deutsche Forschungsgemeinschaft (DFG) through the DFG Cluster of Excellence MATH+ (grant number EXC-2046/1, project ID 390685689).
Appendix A Appendix
A.1 Results from Weak Convexity
Lemma A.1 (Primal gap bound from the quadratic growth).
Let $f$ be a $\rho$-weakly convex function that satisfies the local $\mu$-quadratic growth condition such that $\rho\leq\mu$. Let $x^{*}$ be the unique minimizer of $f$ over $P$ and let $\zeta>0$. Then, the following holds: for all $x\in[f\leq f^{*}+\zeta]\cap P$,
$f(x)-f^{*}\leq\frac{2\mu}{(\mu-\rho)^{2}}\left(\frac{\langle\nabla f(x),x-x^{*}\rangle}{\|x-x^{*}\|}\right)^{2},$
or equivalently,
$\left(\frac{\mu}{2}\right)^{1/2}\left(1-\frac{\rho}{\mu}\right)(f(x)-f^{*})^{1/2}\leq\frac{\langle\nabla f(x),x-x^{*}\rangle}{\|x-x^{*}\|}.$
Proof.
By first applying weak convexity and then the local quadratic growth condition for $x^{*}\in\mathcal{X}^{*}$ with $f(x^{*})=f^{*}$ and for all $x\in[f\leq f^{*}+\zeta]\cap P$, it holds:
$f(x)-f^{*}$ $=f(x)-f(x^{*})$
$\leq\langle\nabla f(x),x-x^{*}\rangle+\frac{\rho}{2}\|x-x^{*}\|^{2}$
$\leq\langle\nabla f(x),x-x^{*}\rangle+\frac{\rho}{\mu}\left(f(x)-f^{*}\right),$
which implies
$\left(1-\frac{\rho}{\mu}\right)\left(f(x)-f^{*}\right)\leq\langle\nabla f(x),x-x^{*}\rangle=\frac{\langle\nabla f(x),x-x^{*}\rangle}{\|x-x^{*}\|}\|x-x^{*}\|.$ (A.1)
Using the local quadratic growth condition again, we have
$\left(1-\frac{\rho}{\mu}\right)\left(f(x)-f^{*}\right)\leq\frac{\langle\nabla f(x),x-x^{*}\rangle}{\|x-x^{*}\|}\left(\frac{2}{\mu}(f(x)-f^{*})\right)^{1/2},$
which provides
$\left(\frac{\mu}{2}\right)^{1/2}\left(1-\frac{\rho}{\mu}\right)(f(x)-f^{*})^{1/2}\leq\frac{\langle\nabla f(x),x-x^{*}\rangle}{\|x-x^{*}\|},$
or equivalently
$f(x)-f^{*}\leq\frac{2\mu}{(\mu-\rho)^{2}}\left(\frac{\langle\nabla f(x),x-x^{*}\rangle}{\|x-x^{*}\|}\right)^{2}.$
Lemma A.2 (Upper bound on primal gap for weakly convex functions).
Suppose that Assumptions 3.1 and 5.2 hold. Let $P$ be a polytope with the pyramidal width $\delta>0$. Let $\mathcal{S}$ denote any set of vertices of $P$ with $x\in\operatorname{conv}\mathcal{S}$. Let $\psi$ be any vector, so that we define $v^{\textup{FW}}=\operatorname*{argmin}_{v\in P}\langle\psi,v\rangle$ and $v^{\textup{A}}=\operatorname*{argmax}_{v\in\mathcal{S}}\langle\psi,v\rangle$. If $\rho/\mu<1$, then it holds that, for all $x\in[f\leq f^{*}+\zeta]\cap P$,
$f(x)-f^{*}\leq\frac{2\mu}{(\mu-\rho)^{2}\delta^{2}}\langle\nabla f(x),v^{\textup{A}}-v^{\textup{FW}}\rangle^{2}.$
Proof.
Using the weak convexity and the local quadratic growth condition of $f$, it holds that, for all $x\in[f\leq f^{*}+\zeta]\cap P$,
$f(x)-f(x^{*})$ $\leq\langle\nabla f(x),x-x^{*}\rangle+\frac{\rho}{2}\|x-x^{*}\|^{2}$
$\leq\langle\nabla f(x),x-x^{*}\rangle+\frac{\rho}{\mu}(f(x)-f(x^{*})),$
which implies
$0\leq\left(1-\frac{\rho}{\mu}\right)(f(x)-f(x^{*}))\leq\langle\nabla f(x),x-x^{*}\rangle,$
where the first inequality follows from $01-\rho/\mu>$ 0. Using (2.10) with $\psi=\nabla f(x)$ and $y=x^{*}$ and the above inequality, we obtain
$\frac{\langle\nabla f(x),v^{\textup{A}}-v^{\textup{FW}}\rangle^{2}}{\delta^{2}}$ $\geq\frac{\langle\nabla f(x),x-x^{*}\rangle^{2}}{\|x-x^{*}\|^{2}}$
$\geq\left(1-\frac{\rho}{\mu}\right)^{2}\frac{(f(x)-f(x^{*}))^{2}}{\|x-x^{*}\|^{2}}$
$\geq\frac{\mu}{2}\left(1-\frac{\rho}{\mu}\right)^{2}\frac{(f(x)-f(x^{*}))^{2}}{f(x)-f(x^{*})}$
$=\frac{\mu}{2}\left(1-\frac{\rho}{\mu}\right)^{2}(f(x)-f(x^{*})),$
where the third inequality holds because of the local quadratic growth condition. ∎
A.2 Local Linear Convergence over Uniformly Convex Sets
In the convex optimization case, Canon and Cullum [Canon1968-sr] established an early lower bound on the convergence rate of the FW algorithm. However, in the special case of $P$ being strongly convex, Garber and Hazan [Garber2015-de] showed that one can improve upon that lower bound. Kerdreux et al. [Kerdreux2021-kc] establish it in the case of $P$ being uniformly convex. We will now carry over this result to establish local linear convergence for the case where $f$ is weakly convex and $L$-smad on $P$.
To this end, we recall the definition of uniformly convex sets.
Definition A.3 ($(\alpha,p)$-uniformly convex set [Braun2022-ew, Definition 2.18][Kerdreux2021-kc, Definition 1.1]).
[…]
$y+\gamma(x-y)+\gamma(1-\gamma)\cdot\alpha\|x-y\|^{p}z\in P.$
Moreover, $P$ is said to be strongly convex if $P$ is $(\alpha,2)$-uniformly convex.
We will use the scaling condition for uniformly convex sets to establish linear convergence.
Proposition A.4 (Scaling inequality [Braun2022-ew, Proposition 2.19][Kerdreux2021-kc, Lemma 2.1]).
Let $P$ be a full dimensional compact $(\alpha,p)$-uniformly convex set, $u$ any non-zero vector, and $v=\operatorname*{argmin}_{y\in P}\langle u,y\rangle$. Then for all $x\in P$
$\frac{\langle u,x-v\rangle}{\|x-v\|^{p}}\geq\alpha\|u\|.$
Now we establish local linear convergence for weakly convex optimization on uniformly convex sets.
Theorem A.5 (Local linear convergence over uniformly convex sets).
Suppose that Assumptions 3.1 and 5.2 with $\phi=\frac{1}{2}\|\cdot\|^{2}$ and $\operatorname{int}\operatorname{dom}\phi=\mathbb{R}^{n}$ hold and that $P$ is $(\alpha,p)$-uniformly convex set. Let $\nabla f$ be bounded away from 0, i.e., $\|\nabla f(x)\|\geq c>0$ for all $x\in P$. Let $D_{\textup{Euc}}:=\sup_{x,y\in P}\|x-y\|$ be the diameter of $P$. Consider the iterates of Algorithm 1 with $\gamma_{t}=\min\left\{\frac{\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle}{L\|x_{t}-v_{t}\|^{2}},1\right\}$. Then, if $\rho<\mu$ and $\rho\leq L$, it holds that
$f(x_{t})-f^{*}\leq\begin{cases}\max\left\{\frac{1}{2}\left(1+\frac{\rho}{\mu}\right),1-\left(1-\frac{\rho}{\mu}\right)\frac{\alpha c}{2L}\right\}^{t-1}LD_{\textup{Euc}}^{2}&\text{if}\ p=2,\\ \left(\frac{1}{2}+\frac{\rho}{2\mu}\right)^{t-1}LD_{\textup{Euc}}^{2}&\text{if}\ 1\leq t\leq t_{0},p\geq 2,\\ \frac{L\left(\left(1-\rho/\mu\right)^{1-p/2}L/\alpha c\right)^{2/(p-2)}}{(1+(1-\rho/\mu)(1/2-1/p)(t-t_{0}))^{p/(p-2)}}=\mathcal{O}(1/t^{p/(p-2)})&\text{if}\ t\geq t_{0},p\geq 2,\end{cases}$
for all $t\geq 1$ where
$t_{0}:=\max\left\{\left\lfloor\log_{\frac{1}{2}\left(1+\frac{\rho}{\mu}\right)}\frac{L\left(\left(1-\rho/\mu\right)^{1-p/2}L/\alpha c\right)^{2/(p-2)}}{LD_{\textup{Euc}}^{2}}\right\rfloor+2,1\right\}.$
Proof.
Let $g_{t}:=\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle$ and $h_{t}:=f(x_{t})-f^{*}$. Lemma 3.3 with $\phi=\frac{1}{2}\|\cdot\|^{2}$ and $\nu=1$ is followed by $f(x_{t})-f(x_{t+1})\geq\frac{g_{t}}{2}\gamma_{t}$. Using Proposition A.4, we have
$h_{t}-h_{t+1}$ $\geq\frac{g_{t}}{2}\min\left\{\frac{g_{t}}{L\|x_{t}-v_{t}\|^{2}},1\right\}$
$\geq\frac{g_{t}}{2}\min\left\{\frac{g_{t}^{1-2/p}\alpha^{2/p}\|\nabla f(x_{t})\|^{2/p}}{L},1\right\}$
$\geq\frac{1}{2}\left(h_{t}-\frac{\rho}{2}\|x-x^{*}\|^{2}\right)\min\left\{\left(h_{t}-\frac{\rho}{2}\|x-x^{*}\|^{2}\right)^{1-2/p}\frac{(\alpha c)^{2/p}}{L},1\right\}$
$\geq\frac{1}{2}\left(1-\frac{\rho}{\mu}\right)\min\left\{\left(1-\frac{\rho}{\mu}\right)^{1-2/p}\frac{(\alpha c)^{2/p}}{L}h_{t}^{1-2/p},1\right\}\cdot h_{t},$
where the third inequality holds from Lemma 5.3 and $\|\nabla f(x)\|\geq c$, and the last inequality holds because of the local quadratic growth property of $f$. The initial bound $h_{1}\leq LD_{\textup{Euc}}^{2}=2LD^{2}$ holds from $\rho\leq L$ and (5.6). For $q=2$, we have $h_{t}-h_{t+1}\geq\frac{1}{2}\left(1-\frac{\rho}{\mu}\right)\min\left\{\frac{\alpha c}{L},1\right\}\cdot h_{t}$, which implies
$h_{t+1}\leq\max\left\{\frac{1}{2}\left(1+\frac{\rho}{\mu}\right),1-\left(1-\frac{\rho}{\mu}\right)\frac{\alpha c}{2L}\right\}\cdot h_{t}.$
Thus, we have the claim. For $p>2$, we use Lemma 2.20 with $c_{0}=LD_{\textup{Euc}}^{2}$, $c_{1}=\frac{1}{2}\left(1-\frac{\rho}{\mu}\right)$, $c_{2}=c_{1}\cdot\left(1-\frac{\rho}{\mu}\right)^{1-2/p}\frac{(\alpha c)^{2/p}}{L}$, and $\theta_{0}=1-2/p$ and obtain the claim. ∎
Note that Assumption 3.1 with $\phi=\frac{1}{2}\|\cdot\|^{2}$ and $\operatorname{int}\operatorname{dom}\phi=\mathbb{R}^{n}$ holds when $f$ is $L$-smooth over $P$. If $\rho\leq L$ does not hold, we can use the initial bound $h_{1}\leq\frac{\rho+L}{2}D_{\textup{Euc}}^{2}$ from (5.6). In that case, a local linear rate in Theorem A.5 is unchanged. Moreover, we have local linear convergence without assuming $\|\nabla f(x)\|>0$.
Theorem A.6 (Local linear convergence over uniformly convex sets without $\|\nabla f(x)\|>0$).
Suppose that Assumptions 3.1 and 5.2 with $\phi=\frac{1}{2}\|\cdot\|^{2}$ and $\operatorname{int}\operatorname{dom}\phi=\mathbb{R}^{n}$ hold and that $P$ is $(\alpha,p)$-uniformly convex set. Let $D_{\textup{Euc}}:=\sup_{x,y\in P}\|x-y\|$ be the diameter of $P$. Consider the iterates of Algorithm 1 with $\gamma_{t}=\min\left\{\frac{\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle}{L\|x_{t}-v_{t}\|^{2}},1\right\}$. Then, if $\rho<\mu$ and $\rho\leq L$, it holds that
$f(x_{t})-f^{*}\leq\begin{cases}\left(\frac{1}{2}+\frac{\rho}{2\mu}\right)^{t-1}LD_{\textup{Euc}}^{2}&\text{if}\ 1\leq t\leq t_{0},p\geq 2,\\ \frac{\left(\left(1-\rho/\mu\right)^{2-p}Lc^{2}/\alpha^{2}\right)^{1/(p-1)}}{(1+(1/2)\cdot(1-\rho/\mu)(1-1/p)(t-t_{0}))^{p/(p-1)}}=\mathcal{O}(1/t^{p/(p-1)})&\text{if}\ t\geq t_{0},p\geq 2,\end{cases}$
for all $t\geq 1$ where $c=\frac{\sqrt{2\mu}}{\mu-\rho}$ and
$t_{0}:=\max\left\{\left\lfloor\log_{\frac{1}{2}\left(1+\frac{\rho}{\mu}\right)}\frac{L\left(\left(1-\rho/\mu\right)^{2-p}Lc^{2}/\alpha^{2}\right)^{1/(p-1)}}{LD_{\textup{Euc}}^{2}}\right\rfloor+2,1\right\}.$
Proof.
Using an argument similar to Theorem A.5, we obtain
$h_{t}-h_{t+1}$ $\geq\frac{g_{t}}{2}\min\left\{\frac{g_{t}}{L\|x_{t}-v_{t}\|^{2}},1\right\}$
$\geq\frac{g_{t}}{2}\min\left\{\frac{g_{t}^{1-2/p}\alpha^{2/p}\|\nabla f(x_{t})\|^{2/p}}{L},1\right\}$
$\geq\frac{1}{2}\left(h_{t}-\frac{\rho}{2}\|x-x^{*}\|^{2}\right)\min\left\{\left(h_{t}-\frac{\rho}{2}\|x-x^{*}\|^{2}\right)^{1-2/p}\frac{\alpha^{2/p}(h_{t}/c^{2})^{1/p}}{L},1\right\}$
$\geq\frac{1}{2}\left(1-\frac{\rho}{\mu}\right)\min\left\{\left(1-\frac{\rho}{\mu}\right)^{1-2/p}\frac{(\alpha c^{-1})^{2/p}}{L}h_{t}^{1-1/p},1\right\}\cdot h_{t},$
where the third inequality holds from the PL inequality (2.8) with $c=\frac{\sqrt{2\mu}}{\mu-\rho}$. Therefore, we use Lemma 2.20 with $c_{0}=LD_{\textup{Euc}}^{2}$, $c_{1}=\frac{1}{2}\left(1-\frac{\rho}{\mu}\right)$, $c_{2}=c_{1}\cdot\left(1-\frac{\rho}{\mu}\right)^{1-2/p}\frac{(\alpha c^{-1})^{2/p}}{L}$, and $\theta_{0}=1-1/p$ and obtain the claim. ∎